### v2

# Polarity transitions induced by symmetry-breaking outer boundary heat flux in rapidly rotating dynamos

Abstract
This study investigates, analytically and numerically, the role of a non-axisymmetric equatorially anti-symmetric lateral variation in heat flux at the outer boundary in polarity transitions in rapidly rotating dynamos. In an unstably stratified fluid, the frequencies of vertical and horizontal (lateral) buoyancy complement each other such that a polarity transition is induced by the boundary anti-symmetry through the suppression of the slow magnetic-Archimedean-Coriolis (MAC) waves at relatively small vertical buoyant forcing. A dipole-dominated dynamo in the low-inertia limit transitions to polarity-reversing and multipolar states in succession as the relative intensity of horizontal buoyancy is progressively increased for a fixed vertical buoyancy. In the same parameter space, an equatorially symmetric heat flux variation does not induce a polarity transition. A composite boundary heterogeneity consisting of comparable magnitudes of symmetric and anti-symmetric variations induces the polarity transition at a horizontal buoyancy of the same order as that for the transition induced by a purely anti-symmetric variation. This makes the analysis relevant to Earth’s core, which convects in response to a composite heat flux variation in the lowermost mantle. A heterogeneity with a dominant equatorially symmetric component does not favour polarity transitions, likely producing long periods in Earth’s history without reversals. The fact that compositional buoyancy is much stronger than thermal buoyancy, together with the known order of magnitude of the peak field intensity in the inertia-free limit, indicates a large lower-mantle heat flux heterogeneity of $O(10)$ times the mean superadiabatic heat flux at the core–mantle boundary.
1 Introduction
It is well-known that the Earth’s magnetic field is generated by thermochemical convection in the outer core. Compositional convection arises from the continual release of light elements that accompanies the progressive growth of the inner core. Thermal convection occurs primarily by secular cooling of the core as the planet cools in time. Mantle convection produces a laterally heterogeneous heat flux at the core–mantle boundary (CMB), which changes in geological time (Olson et al., 2010). Since the convective turnover time in the lower mantle is significantly longer than in the outer core (Zhang and Gubbins, 1992; Holme et al., 2015), these lateral variations create a quasi-stationary boundary condition that shapes outer core convection.
Based on several observations, it has been argued that the lower mantle may influence the geodynamo. The persistence of high-latitude flux patch concentrations in the paleomagnetic time average (Gubbins and Kelly, 1993; Carlut and Courtillot, 1998; Johnson et al., 2003), the preferred paths of virtual geomagnetic poles during polarity transitions (Laj et al., 1991; Love, 2000), and the low secular variation in the Pacific (Fisk, 1931; Doell and Cox, 1971) are examples. Paleomagnetic data also suggest that the frequency of geomagnetic reversals changes on the time scale of mantle convection (Jones, 1977; McFadden and Merrill, 1984; Larson and Olson, 1991; Lhuillier et al., 2013). In addition, recent laboratory experiments (Sahoo and Sreenivasan, 2020) indicate that a large lower-mantle heterogeneity gives rise to an east–west dichotomy in core convection, which in turn explains the relative instability of the high-latitude magnetic flux lobes in the Western hemisphere (Jackson et al., 2000).
The first self-consistent dynamo models with inhomogeneous CMB heat flux (Glatzmaier et al., 1999) showed that the frequency of geomagnetic reversals depends on the heat flux pattern. This CMB heat flux variation was estimated from the shear wave velocity anomalies obtained from seismic tomography, under the assumption that shear wave velocity is determined by temperature and not by composition (Su et al., 1994; Masters et al., 1996). However, compositional heterogeneity can also cause significant seismic wave velocity variations in the lower mantle (Koelemeijer et al., 2012; Gülcher et al., 2021). Several numerical dynamo models (Olson and Christensen, 2002; Christensen and Olson, 2003; Aubert et al., 2007; Sreenivasan and Gubbins, 2011; Olson and Amit, 2014; Mound and Davies, 2023) have used CMB heat flux variations that linearly correlate with seismic shear wave velocity anomalies. Additionally, various single harmonics have been used to understand their effect in isolation on convection and the magnetic field (Kutzner and Christensen, 2004; Takahashi et al., 2008; Coe and Glatzmaier, 2006; Courtillot and Olson, 2007; Stanley et al., 2008; Amit et al., 2011). It has been shown that both an enhanced equatorial heat flux (Glatzmaier et al., 1999; Olson et al., 2010) and a North–South hemispherical heat flux variation (Frasson et al., 2025) can produce reversals in numerical dynamos. Mantle convection models indicate that the dominant spherical harmonic in the mantle flow switches between non-axisymmetric $l=1$ and $l=2$ (Zhong et al., 2007; Yoshida, 2008; Olson et al., 2010), where $l$ is spherical harmonic degree. During supercontinent aggregation, the non-axisymmetric $l=1$ mantle flow tends to dominate, whereas $l=2$ becomes predominant after supercontinent breakup. Therefore, it makes sense to focus on the effect of non-axisymmetric degree 1 and 2 heat flux patterns at the outer boundary. Of these, equatorially anti-symmetric conditions are likely to trigger polarity transitions (e.g. Takahashi et al., 2008). The equatorially anti-symmetric part of the lower-mantle heat flux can become significant for large lateral variations in boundary heat flux and thus influence the polarity of the dynamo. The present study analyses this problem in the rapidly rotating, strongly driven regime of a planetary dynamo.
A heterogeneous heat flux pattern at the outer boundary induces a steady mean flow, termed the thermal wind. A study of the onset of convection in a plane layer in the presence of a thermal wind (Teed et al., 2010) suggests that dynamo action might be possible even in subadiabatic conditions. Near convective onset, an equatorially symmetric boundary heterogeneity in heat flux likely favours dynamo action at large magnetic Prandtl number whereas an equatorially anti-symmetric heterogeneity can induce polarity reversals (Sahoo et al., 2016). In a thermally driven dynamo, an equatorially anti-symmetric heat flux heterogeneity induces a mean axial temperature gradient at the equator, which modifies the buoyancy profile of the homogeneous state, and hence the buoyancy frequency. For a sufficiently large anti-symmetric heterogeneity, the resultant buoyancy frequency $\omega_{A}$ matches the Alfvén wave frequency $\omega_{M}$, causing the suppression of the slow Magnetic–Archimedean–Coriolis (MAC), or magnetostrophic, waves in the dynamo. In a recent study (Majumder et al., 2024), variations in the mean outer boundary heat flux were interpreted as changes in the Rayleigh number, and the condition $|\omega_{A}|\approx|\omega_{M}|$ was shown to be associated with the loss of kinetic helicity from these waves, in turn leading to the collapse of the axial dipole magnetic field.
In an unstably stratified fluid, convection occurs when a superadiabatic gradient is present in the basic state. Within this medium, isolated density perturbations give rise to fast and slow MAC waves whose frequencies are obtainable in the Boussinesq limit (Braginsky, 1967; Busse et al., 2007; Majumder et al., 2024). In a rapidly rotating fluid where the magnitude of the inertial wave frequency $|\omega_{C}|\gg|\omega_{M}|,\,|\omega_{A}|$, the real part of the slow wave frequency is approximated by (e.g. Majumder et al., 2024),
$\omega_{s}\approx\dfrac{\omega_{M}^{2}}{\omega_{C}}\left(1+\dfrac{\omega_{A}^{2}}{\omega_{M}^{2}}\right)^{1/2}.$ (1)
The spontaneous generation of slow MAC waves in a convection-driven dynamo that evolves from a small seed magnetic field (figure S1, Supplementary material) indicates that these waves exist in an unstably stratified fluid where $\omega_{A}^{2}<0$, $\omega_{M}^{2}>0$ and $|\omega_{M}|\geq|\omega_{A}|$ in (1). As the dynamo passes through a chaotic multipolar state, the slow waves are excited by localized balances between the Lorentz, buoyancy and Coriolis forces, where the Lorentz force is made up of the non-dipolar field (figure S2, Supplementary Material). The helicity of the slow waves, which is at least as high as that of the fast waves, is essential for the formation of the dipole field from this multipolar state (Varma and Sreenivasan, 2022). The hydrodynamic dynamo, where the Lorentz force is absent, does not produce the dipole from a seed field (Sreenivasan and Kar, 2018) since the slow MAC waves are not generated in the first place. As the buoyant forcing is progressively increased in the dipolar regime, $|\omega_{M}|$ attains its highest value, upon which a state where $|\omega_{A}|\approx|\omega_{M}|$ ensues. Here, (1) predicts the disappearance of the slow MAC waves, which brings about the collapse of the axial dipole field (Majumder et al., 2024).
The present study builds on previous work, with a focus on examining how heterogeneity in the outer boundary heat flux causes the dipole–multipole transition. Since polarity transitions occur when $|\omega_{A}|\approx|\omega_{M}|$, and the resultant buoyancy frequency $\omega_{A}$ is made up of both vertical (radial) and horizontal (lateral) parts, it is in principle possible for a weak vertical buoyancy to be compensated by a strong horizontal buoyancy in order to match $|\omega_{M}|$. This complementarity between vertical and horizontal buoyancy suggests that polarity transitions can arise even under weakly driven thermal convection, provided there is a sufficiently large heat flux heterogeneity. For a fixed vertical buoyancy, polarity reversals may exist in a narrow range of horizontal buoyancy states that lie between the dipolar and multipolar regimes. As in dynamos with uniform boundary heat flux (Majumder et al., 2024), we anticipate that the dipole–multipole transition would be self-similar, in that the resultant Rayleigh number based on the characteristic energy-containing length scale bears the same linear relationship with the square of the peak magnetic field at the transition, regardless of the scale of energy injection.
The formation and growth of the inner core make the primary source of buoyancy for outer core convection compositional, which is thought to contribute up to 80% of the total convective power (Lister and Buffett, 1995). The weaker source of buoyancy, derived from thermal convection, is influenced by heat flux heterogeneity at the CMB. In two-component convection, the vertical buoyancies of composition and temperature, together with the horizontal buoyancy from the lateral heat flux variation, make up the resultant buoyant forcing. Using this complementarity of buoyancies and the condition for vanishing slow MAC waves, an upper bound for the lower-mantle heterogeneity that would admit an axial dipole field may be obtained.
In §2, we consider the evolution of an isolated density disturbance in an unstably stratified fluid subject to a uniform magnetic field, background rotation and a lateral temperature variation. Here, the gravity and magnetic field axes coincide in Cartesian geometry, which is referred to as the ‘equatorial radial configuration’ by Loper et al. (2003). This analysis quantifies the resultant buoyancy based on the vertical and horizontal buoyancies and shows their complementary role in the evolution of fast and slow MAC waves. The Cartesian magnetoconvection model with non-zero mean axial temperature gradient serves as the basis for the study on the effect of equatorially anti-symmetric, or symmetry-breaking, outer boundary heat flux on wave motions in rapidly rotating spherical dynamos, described in §3. In a dynamo where the nonlinear inertial forces are small, an equatorially anti-symmetric condition causes polarity transitions not by breaking the symmetry of the columnar vortices but rather by selectively suppressing the slow MAC waves that already exist in the dipolar dynamo. This study also considers composite boundary heat flux patterns obtained by adding equatorially symmetric and anti-symmetric patterns in a known proportion in order to extend the analysis to lower-mantle heat flux anomalies. §4 analyses two-component linear magnetoconvection and suggests an upper bound for the horizontal buoyancy that would admit a dipolar dynamo for a given thermal power ratio. In §5, we discuss the implications of our results for Earth and propose future work.
2 Evolution of a density disturbance under rapid rotation, lateral temperature variation and a magnetic field
2.1 Problem set-up and governing equations
Figure 1: The initial state of a density perturbation $\rho^{\prime}$ that evolves in an unstably stratified fluid subject to a uniform magnetic field $\bm{B}_{0}=B_{0}\hat{\bm{e}}_{y}$, background rotation $\bm{\varOmega}=\varOmega\hat{\bm{e}}_{z}$, and gravity $\bm{g}=-g\hat{\bm{e}}_{y}$ in Cartesian coordinates $(x,y,z)$. The lateral variation in temperature produces a mean flow $\bm{u}_{0}=u_{c}z\,\hat{\bm{e}}_{x}$.
A localized density perturbation $\rho^{\prime}$ is situated in an unstably stratified fluid subject to a uniform magnetic field and rapid rotation. Since $\rho^{\prime}$ is related to a temperature perturbation $\Theta$ by $\rho^{\prime}=-\rho\alpha\Theta$, where $\rho$ is the ambient density and $\alpha$ is the coefficient of thermal expansion, an initial temperature perturbation is chosen in the form
$\Theta_{0}=C\,\exp\left[-(x^{2}+y^{2}+z^{2})/\delta^{2}\right],$ (2)
where $C$ is a constant and $\delta$ is the length scale of the perturbation. Figure 1 shows the initial perturbation, which evolves under the influence of gravity $\bm{g}=-g\hat{\bm{e}}_{y}$, background rotation $\bm{\varOmega}=\varOmega\hat{\bm{e}}_{z}$, and a uniform magnetic field $\bm{B}_{0}=B_{0}\hat{\bm{e}}_{y}$ in Cartesian coordinates $(x,y,z)$. The basic state temperature $T_{0}$ is assumed to have constant variations in both the vertical ($y$) and a horizontal ($z$) direction (Teed et al., 2010), which results in a steady mean flow $\bm{u}_{0}$ satisfying the thermal wind equation
$2\varOmega\frac{\partial\bm{u}_{0}}{\partial z}=-g\alpha\hat{\bm{e}}_{y}\times\nabla T_{0}.$ (3)
Here, the mean flow corresponds to a constant shear,
$\bm{u}_{0}=u_{c}z\,\hat{\bm{e}}_{x},$ (4)
which on substitution in (3) in scalar form,
$2\varOmega\,u_{c}=-g\alpha\dfrac{\partial T_{0}}{\partial z},$ (5)
gives the basic state temperature,
$T_{0}=\beta_{y}y+\beta_{z}z,$ (6)
where $\beta_{y}=\partial T_{0}/\partial y<0$ is the mean vertical temperature gradient and $\beta_{z}=-2\varOmega u_{c}/g\alpha>0$ is the mean horizontal (lateral) temperature gradient.
The initial temperature perturbation (2) gives rise to a velocity field $\bm{u}$, which interacts with the ambient magnetic field $\bm{B}_{0}$ to produce the induced magnetic field $\bm{b}$. The initial velocity perturbation and induced field are both zero. Since all the variables are decomposed into their mean and perturbation parts, e.g.
$T=T_{0}+\Theta,\quad\bm{U}=\bm{u}_{0}+\bm{u},\quad\bm{B}=\bm{B}_{0}+\bm{b},$ (7a–c)
the following linearized equations give the evolution of $\bm{u}$, $\bm{b}$ and $\Theta$:
$\frac{\partial\bm{u}}{\partial t}+u_{c}\,z\frac{\partial\bm{u}}{\partial x}+u_{c}u_{z}\hat{\bm{e}}_{x}+2\varOmega\hat{\bm{e}}_{z}\times\bm{u}=-\frac{1}{\rho}\nabla p^{\star}+\frac{1}{\rho}\bm{j}\times\bm{B}_{0}+g\alpha\Theta\hat{\bm{e}}_{y}+\nu\nabla^{2}\bm{u},$ (8)
$\frac{\partial\bm{b}}{\partial t}+u_{c}\,z\frac{\partial\bm{b}}{\partial x}=u_{c}b_{z}\hat{\bm{e}}_{x}+(\bm{B}_{0}\cdot\nabla)\bm{u}+\eta\nabla^{2}\bm{b},$ (9)
$\frac{\partial\Theta}{\partial t}+u_{c}z\frac{\partial\Theta}{\partial x}+u_{y}\beta_{y}+u_{z}\beta_{z}=\kappa\nabla^{2}\Theta,$ (10)
$\nabla\cdot\bm{u}=0,$ (11)
$\nabla\cdot\bm{b}=0,$ (12)
where $\bm{j}$ is the electric current density, $\nu$ is the kinematic viscosity, $\kappa$ is the thermal diffusivity, $\eta$ is the magnetic diffusivity and $p^{*}=p-(\rho/2)|{\bm{\Omega}}\times{\bm{x}}|^{2}$ is a modified pressure.
2.2 Solutions for the perturbation velocity field
Taking the curl of equation (8) and (9) and eliminating $\bm{j}$, we obtain
$P_{\nu}$ $P_{\eta}\bm{\zeta}=2(\bm{\Omega}\cdot\nabla)P_{\eta}\bm{u}$ (13)
$-u_{c}P_{\eta}\bigg[-\frac{\partial u_{y}}{\partial x}\hat{\bm{e}}_{x}+\bigg(\frac{\partial u_{x}}{\partial x}+\frac{\partial u_{z}}{\partial z}\bigg)\hat{\bm{e}}_{y}-\frac{\partial u_{z}}{\partial y}\hat{\bm{e}}_{z}\bigg]+g\alpha P_{\eta}\bigg(-\frac{\partial\Theta}{\partial z}\hat{\bm{e}}_{x}+\frac{\partial\Theta}{\partial x}\hat{\bm{e}}_{z}\bigg)$
$+u_{c}\frac{(\bm{B}_{0}\cdot\nabla)}{\rho\mu}\bigg[\frac{\partial b_{y}}{\partial x}\hat{\bm{e}}_{x}+\bigg(-\frac{\partial b_{x}}{\partial x}+\frac{\partial b_{z}}{\partial z}\bigg)\hat{\bm{e}}_{y}-\frac{\partial b_{z}}{\partial y}\hat{\bm{e}}_{z}\bigg]+\frac{(\bm{B}_{0}\cdot\nabla)^{2}}{\rho\mu}\bm{\zeta}.$
where $P_{\nu}=\frac{\partial}{\partial t}+u_{c}z\frac{\partial}{\partial x}-\nu\nabla^{2}$, $P_{\eta}=\frac{\partial}{\partial t}+u_{c}z\frac{\partial}{\partial x}-\eta\nabla^{2}$, $\bm{\zeta}=\nabla\times\bm{u}$ is the vorticity and $\mu$ is the magnetic permeability. Successive elimination of $\bm{\zeta}$ and $\Theta$ through the $z$ component of the curl of (13), the $z$ component of (9) and (10) gives
$P^{2}P_{\kappa}(-\nabla^{2}u_{z})=P\bigg[-g\alpha P_{\eta}\frac{\partial^{2}}{\partial y\partial z}(\beta_{y}u_{y}+\beta_{z}u_{z})+u_{c}P_{\kappa}\frac{B_{0}}{\rho\mu}\frac{\partial}{\partial y}\bigg(-\frac{\partial^{2}b_{x}}{\partial x^{2}}+\frac{\partial^{2}b_{z}}{\partial x\partial z}-\frac{\partial^{2}b_{y}}{\partial x\partial y}\bigg)\bigg]$ (14)
$\hskip 20.00003pt\ignorespaces+2\varOmega\frac{\partial}{\partial z}P_{\kappa}\bigg[2\varOmega\frac{\partial}{\partial z}P_{\eta}^{2}u_{z}+u_{c}P_{\eta}^{2}\frac{\partial u_{z}}{\partial y}-u_{c}\frac{B_{0}^{2}}{\rho\mu}\frac{\partial^{3}u_{z}}{\partial y^{3}}\bigg]-2\varOmega g\alpha\frac{\partial^{2}}{\partial x\partial z}P_{\eta}^{2}(\beta_{y}u_{y}+\beta_{z}u_{z}),$
where $P=P_{\nu}P_{\eta}-B_{0}^{2}/\rho\mu$ and $P_{\kappa}=\frac{\partial}{\partial t}+u_{c}z\frac{\partial}{\partial x}-\kappa\nabla^{2}$. For zero horizontal variation in temperature, $\beta_{z},\,u_{c}\to 0$, and (14) reduces to the classical MAC wave equation (Braginsky, 1967; Busse et al., 2007, pp. 165–168). On the other hand, since $x$ is not a preferred horizontal direction, setting $\partial(\,\,)/\partial x=0$ in (14) (e.g. Hathaway et al., 1979) gives
$\bigg[\bigg(\frac{\partial}{\partial t}-\nu\nabla_{*}^{2}\bigg)\bigg(\frac{\partial}{\partial t}-\eta\nabla_{*}^{2}\bigg)-\frac{B_{0}^{2}}{\rho\mu}\,\frac{\partial^{2}}{\partial y^{2}}\bigg]^{2}\bigg(\frac{\partial}{\partial t}-\kappa\nabla_{*}^{2}\bigg)(-\nabla_{*}^{2}\hat{u}_{z})$ (15)
$=-g\alpha\bigg(\frac{\partial}{\partial t}-\eta\nabla_{*}^{2}\bigg)\bigg[\bigg(\frac{\partial}{\partial t}-\nu\nabla_{*}^{2}\bigg)\bigg(\frac{\partial}{\partial t}-\eta\nabla_{*}^{2}\bigg)-\frac{B_{0}^{2}}{\rho\mu}\,\frac{\partial^{2}}{\partial y^{2}}\bigg]\,\frac{\partial^{2}}{\partial y\partial z}(\beta_{y}\hat{u}_{y}+\beta_{z}\hat{u}_{z})$
$+2\varOmega\,\frac{\partial}{\partial z}\bigg(\frac{\partial}{\partial t}-\kappa\nabla_{*}^{2}\bigg)\bigg[2\varOmega\frac{\partial}{\partial z}\bigg(\frac{\partial}{\partial t}-\eta\nabla_{*}^{2}\bigg)^{2}+u_{c}\frac{\partial}{\partial y}\bigg(\frac{\partial}{\partial t}-\eta\nabla_{*}^{2}\bigg)^{2}-u_{c}\frac{B_{0}^{2}}{\rho\mu}\frac{\partial^{3}}{\partial y^{3}}\bigg]\hat{u}_{z},$
where $\nabla_{*}^{2}=\partial^{2}/\partial y^{2}+\partial^{2}/\partial z^{2}$.
Now, applying the two dimensional Fourier transform defined by
$\mathcal{F}\big(\mathbf{A}\big)=\hat{\mathbf{A}}=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\mathbf{A}\mbox{e}^{-\mathrm{i}\,(k_{y}\,y+k_{z}\,z)}\,\mathrm{d}y\,\mathrm{d}z,$ (16)
to (15), the following homogeneous equation is obtained:
$\bigg[\bigg(\frac{\partial}{\partial t}$ $+\nu k^{2}\bigg)\bigg(\frac{\partial}{\partial t}+\eta k^{2}\bigg)+\frac{B_{0}^{2}\,k_{y}^{2}}{\rho\mu}\bigg]^{2}\bigg(\frac{\partial}{\partial t}+\kappa k^{2}\bigg)k^{2}\,\hat{u}_{z}$ (17)
$=$ $-g\alpha k_{z}^{2}\bigg(\frac{\partial}{\partial t}+\eta k^{2}\bigg)\bigg[\bigg(\frac{\partial}{\partial t}+\nu k^{2}\bigg)\bigg(\frac{\partial}{\partial t}+\eta k^{2}\bigg)+\frac{B_{0}^{2}\,k_{y}^{2}}{\rho\mu}\bigg]\left(\beta_{y}-\beta_{z}\frac{k_{y}}{k_{z}}\right)\,\hat{u}_{z}$
$+2\varOmega\,\mathrm{i}\,k_{z}\bigg(\frac{\partial}{\partial t}+\kappa k^{2}\bigg)\bigg[2\varOmega\,\mathrm{i}\,k_{z}\bigg(\frac{\partial}{\partial t}+\eta k^{2}\bigg)^{2}+u_{c}\mathrm{i}k_{y}\bigg(\frac{\partial}{\partial t}+\eta k^{2}\bigg)^{2}+u_{c}\mathrm{i}k_{y}\frac{B_{0}^{2}\,k_{y}^{2}}{\rho\mu}\bigg]\,\hat{u}_{z},$
where $k^{2}=k_{y}^{2}+k_{z}^{2}$ and the following substitution is made from the continuity equation (11):
$\hat{u}_{y}=-\frac{k_{z}}{k_{y}}\hat{u}_{z}.$ (18)
Seeking a plane wave solution of the form $\hat{u}_{z}\sim\mbox{e}^{\mathrm{i}\lambda t}$, we obtain,
$\big[\big(\mathrm{i}\lambda+\omega_{\nu}\big)$ $\big(\mathrm{i}\lambda+\omega_{\eta}\big)+\omega_{M}^{2}\big]^{2}\big(\mathrm{i}\lambda+\omega_{\kappa}\big)=-\omega_{A,V}^{2}\big(\mathrm{i}\lambda+\omega_{\eta}\big)\big[\big(\mathrm{i}\lambda+\omega_{\nu}\big)\big(\mathrm{i}\lambda+\omega_{\eta}\big)+\omega_{M}^{2}\big]$ (19)
$+\omega_{A,H}^{2}\big(\mathrm{i}\lambda+\omega_{\eta}\big)\big[\big(\mathrm{i}\lambda+\omega_{\nu}\big)\big(\mathrm{i}\lambda+\omega_{\eta}\big)+\omega_{M}^{2}\big]-\omega_{C}^{2}\big(\mathrm{i}\lambda+\omega_{\kappa}\big)\big(\mathrm{i}\lambda+\omega_{\eta}\big)^{2}$
$+\omega_{A,H}^{2}\big(\mathrm{i}\lambda+\omega_{\kappa}\big)\big[\big(\mathrm{i}\lambda+\omega_{\eta}\big)^{2}+\omega_{M}^{2}\big],$
where
$\begin{aligned} &\omega_{C}^{2}=4\varOmega^{2}k_{z}^{2}/k^{2},\quad\omega_{A,V}^{2}=g\alpha\beta_{y}k_{z}^{2}/k^{2},\quad\omega_{A,H}^{2}=g\alpha\beta_{z}k_{y}k_{z}/k^{2},\\ &\omega_{M}^{2}=B_{0}^{2}\,k_{y}^{2}/\rho\mu=V_{M}^{2}k_{y}^{2},\quad\omega_{\eta}^{2}=\eta^{2}k^{4},\quad\omega_{\kappa}^{2}=\kappa^{2}k^{4},\quad\omega_{\nu}^{2}=\nu^{2}k^{4},\end{aligned}$ (20a–g)
represent the squares of the frequencies of linear inertial waves, vertical and horizontal parts of the buoyancy, Alfvén waves, magnetic diffusion, thermal diffusion and viscous diffusion, respectively. Here, $V_{M}=B_{0}/\sqrt{\rho\mu}$ is the Alfvén wave velocity. In line with a recent study (Sreenivasan and Maurya, 2021) where both viscous and thermal diffusion are much smaller than magnetic diffusion, the characteristic equation has the following form in the limit of $\omega_{\kappa},\omega_{\nu}\to 0$:
$\lambda^{5}-2\mathrm{i}\omega_{\eta}\lambda^{4}$ $-(\omega_{C}^{2}+\omega_{\eta}^{2}+2\omega_{M}^{2}+\omega_{A,V}^{2}-2\omega_{A,H}^{2})\lambda^{3}+2\mathrm{i}\omega_{\eta}(\omega_{C}^{2}+\omega_{M}^{2}+\omega_{A,V}^{2}-2\omega_{A,H}^{2})\lambda^{2}$ (21)
$+(\omega_{C}^{2}\omega_{\eta}^{2}+\omega_{M}^{4}+(\omega_{M}^{2}+\omega_{\eta}^{2})(\omega_{A,V}^{2}-2\omega_{A,H}^{2}))\lambda-\mathrm{i}\omega_{\eta}\omega_{M}^{2}(\omega_{A,V}^{2}-\omega_{A,H}^{2})=0.$
The general solution for $\hat{u}_{z}$ is given by
$\hat{u}_{z}=\sum_{m=1}^{5}D_{m}\mbox{e}^{\mathrm{i}\lambda_{m}t},$ (22)
where the coefficients $D_{m}$ are evaluated from the initial conditions of $\hat{u}_{z}$ and its derivatives ($\S$2.3 below). Of the five terms in the expansion on the right-hand side of equation (22), two terms represent oppositely travelling fast MAC waves, two other terms represent oppositely travelling slow MAC waves, and the fifth term represents the overall growth of the velocity perturbation.
2.3 Evaluation of spectral coefficients
From (22), the initial conditions for $\hat{u}_{z}$ and its time derivatives are given by
$\mathrm{i}^{n}\sum_{m=1}^{5}D_{m}\lambda_{m}^{n}=\bigg(\frac{\partial^{n}\hat{u}_{z}}{\partial t^{n}}\bigg)_{t=0}=a_{n+1},\ n=0,1,2,3,4.$ (23)
Algebraic simplifications give the right-hand sides of (23) as follows:
$a_{1}$ $=\hat{u}_{z}|_{t=0}=0,$ (24)
$a_{2}$ $=\frac{\partial{\hat{u}_{z}}}{\partial{t}}|_{t=0}=-g\alpha\frac{k_{y}k_{z}}{k^{2}}\hat{\Theta}_{0},$
$a_{3}$ $=\frac{\partial^{2}\hat{u}_{z}}{\partial{t^{2}}}|_{t=0}=0,$
$a_{4}$ $=\frac{\partial^{3}{\hat{u}_{z}}}{\partial{t^{3}}}|_{t=0}=-(\omega_{M}^{2}+\omega_{C}^{2}+\omega_{A,V}^{2}-2\omega_{A,H}^{2})~a_{2},$
$a_{5}$ $=\frac{\partial^{4}{\hat{u}_{z}}}{\partial{t^{4}}}|_{t=0}=\omega_{M}^{2}\omega_{\eta}~a_{2}.$
The coefficients $D_{m}$ may now be obtained using the roots of equation (21). For example, we obtain,
$D_{1}$ $=\dfrac{a_{5}-\mathrm{i}\,a_{4}(\lambda_{2}+\lambda_{3}+\lambda_{4}+\lambda_{5})+\mathrm{i}\,a_{2}(\lambda_{2}\lambda_{4}\lambda_{5}+\lambda_{3}\lambda_{4}\lambda_{5}+\lambda_{2}\lambda_{3}\lambda_{4}+\lambda_{2}\lambda_{3}\lambda_{5})}{(\lambda_{1}-\lambda_{2})(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{4})(\lambda_{1}-\lambda_{5})},$ (25)
$D_{3}$ $=\dfrac{a_{5}-\mathrm{i}\,a_{4}(\lambda_{1}+\lambda_{2}+\lambda_{4}+\lambda_{5})+\mathrm{i}\,a_{2}(\lambda_{1}\lambda_{4}\lambda_{5}+\lambda_{2}\lambda_{4}\lambda_{5}+\lambda_{1}\lambda_{2}\lambda_{4}+\lambda_{1}\lambda_{2}\lambda_{5})}{(\lambda_{3}-\lambda_{1})(\lambda_{3}-\lambda_{2})(\lambda_{3}-\lambda_{4})(\lambda_{3}-\lambda_{5})},$ (26)
for the forward-travelling fast and slow wave solutions respectively. We separate the fast and slow MAC wave parts of the general solution, which is a linear superposition of the two wave solutions (Sreenivasan and Maurya, 2021). For example,
$\hat{u}_{z,f}$ $=D_{1}\mathrm{e}^{\mathrm{i}{\lambda}_{1}t}+D_{2}\mathrm{e}^{\mathrm{i}{\lambda}_{2}t},$ (27)
$\hat{u}_{z,s}$ $=D_{3}\mathrm{e}^{\mathrm{i}{\lambda}_{3}t}+D_{4}\mathrm{e}^{\mathrm{i}{\lambda}_{4}t},$
where the subscripts $f$ and $s$ in the left-hand sides of (27) denote the fast and slow wave parts of the solution.
2.4 Complementarity of vertical and horizontal buoyancies
For the rapidly rotating regime given by $|\omega_{C}|\gg|\omega_{M}|\gg|\omega_{A,V}|,|\omega_{A,H}|\gg|\omega_{\eta}|$, the roots of the characteristic equation (21) are approximated by
$\lambda_{1,2}$ $\approx$ $\pm\bigg(\omega_{C}+\frac{\omega_{M}^{2}}{\omega_{C}}\bigg)+\mathrm{i}\,\frac{\omega_{M}^{2}\omega_{\eta}}{\omega_{C}^{2}},$ (28)
$\lambda_{3,4}$ $\approx$ $\pm\bigg(\frac{\omega_{M}^{2}}{\omega_{C}}+\frac{\omega_{A,V}^{2}-2\omega_{A,H}^{2}}{2\omega_{C}}\bigg)+\mathrm{i}\,\omega_{\eta}\,\bigg(1-\frac{\omega_{A,V}^{2}-\omega_{A,H}^{2}}{2\omega_{M}^{2}}\bigg),$ (29)
$\lambda_{5}$ $\approx$ $\mathrm{i}\,\frac{\omega_{\eta}(\omega_{A,V}^{2}-\omega_{A,H}^{2})}{\omega_{M}^{2}},$ (30)
following the procedure in Sreenivasan and Maurya (2021). Here, the slow MAC waves of frequency $\lambda_{3,4}$ are damped on the time scale $(\omega_{\eta})_{0}^{-1}=(\eta k_{0}^{2})^{-1}\sim\delta^{2}/\eta$, where the subscript ‘0’ represents the initial state of the buoyancy disturbance (see also Sreenivasan and Maurya, 2021).
From (29), the resultant buoyancy frequency is given by
${\omega_{A}}^{2}=$ $\omega_{A,V}^{2}-2\omega_{A,H}^{2}=\omega_{A,V}^{2}\bigg(1-2\,\beta^{\star}\frac{k_{y}}{k_{z}}\bigg),$ (31)
where $\beta^{\star}=\beta_{z}/\beta_{y}$. If $\beta^{\star}=0$, the approximate classical MAC wave roots are recovered. In the limit of $|\omega_{C}|\gg|\omega_{M}|$, $|\omega_{A}|$, the real parts of $\lambda_{3,4}$ may be expressed as (Braginsky, 1967),
$\mbox{Re}(\lambda_{3,4})\approx\pm\frac{\omega_{M}^{2}}{\omega_{C}}\,\bigg(1+\frac{{\omega_{A}}^{2}}{\omega_{M}^{2}}\bigg)^{1/2},$ (32)
which implies that slow MAC waves would vanish in an unstably stratified fluid with ${\omega_{A}}^{2}<0$ as $|\omega_{A}|$ nears $|\omega_{M}|$. Since $\beta_{y}<0$ and $\beta_{z}>0$ in equation (6), $\omega_{A,V}^{2}<0$ and $\omega_{A,H}^{2}>0$ in equation (31). These two buoyancy frequencies contribute to the resultant buoyancy frequency. An increase in either $|\omega_{A,H}|$ or $|\omega_{A,V}|$ results in an increase in $|\omega_{A}|$, which may match $|\omega_{M}|$ and thereby suppress the slow MAC waves. While a large vertical buoyancy requires a small horizontal buoyancy to suppress these waves, a small vertical buoyancy together with a large horizontal buoyancy would produce the same result. As we shall see later in §4, the complementarity of the buoyancy frequencies places a bound on the magnitude of the horizontal buoyancy in two-component convection.
2.5 Fast and slow MAC waves under unstable stratification and a thermal wind
The evolution of velocity and the induced magnetic field is obtained from the solution of the initial value problem. The solution to the problem is obtained for times much shorter than the time scale for the exponential increase of the perturbations. The Lehnert number $Le$ and the magnetic Ekman number $E_{\eta}$ based on the length scale of the perturbation are used to describe the parameter regime,
$Le=\dfrac{V_{M}}{2\varOmega\delta}\approx\bigg(\frac{\omega_{M}}{\omega_{C}}\bigg)_{\!0},\quad E_{\eta}=\dfrac{\eta}{2\varOmega\delta^{2}}\approx\bigg(\frac{\omega_{\eta}}{\omega_{C}}\bigg)_{\!0},$ (33a,b)
The relative intensity of horizontal buoyancy is measured by the ratio $Ra_{\ell,H}/Ra_{\ell}$, where $Ra_{\ell}$ is the resultant local Rayleigh number given by the sum of the local vertical and horizontal Rayleigh numbers,
$Ra_{\ell,V}=\dfrac{g\alpha|\beta_{y}|\delta^{2}}{2\varOmega\eta},\quad Ra_{\ell,H}=2\dfrac{g\alpha|\beta_{z}|\delta^{2}}{2\varOmega\eta}\frac{k_{y}}{k_{z}},$ (34a,b)
so that
$Ra_{\ell}=\dfrac{g\alpha|\beta|\delta^{2}}{2\varOmega\eta},$ (35)
based on the resultant temperature gradient,
$\beta=\beta_{y}-2\beta_{z}k_{y}/k_{z}.$ (36)
(a)            (b)            (c)
Refer to caption Refer to caption Refer to caption
Figure 2: Evolution of $u_{z}^{2}$ on the $y$-$z$ plane at $x=0$ with time (measured in units of the magnetic diffusion time $t_{\eta}$) for $Le=0.03$ and $E_{\eta}=2\times 10^{-5}$. The snapshots are at (a) $t/t_{\eta}=1\times 10^{-4}$, (b) $t/t_{\eta}=1\times 10^{-3}$ and (c) $t/t_{\eta}=1\times 10^{-2}$ for $Ra_{\ell,{H}}/Ra_{\ell}=0.4$ and $|\omega_{A}^{2}/\omega_{M}^{2}|=0.1$.
Figure 2 shows the evolution of the axial kinetic energy density of the perturbation. Here, the real-space $z$ velocity is obtained from the inverse Fourier transform of (22),
$u_{z}=\frac{1}{(2\pi)^{2}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\hat{u}_{z}\mbox{e}^{\mathrm{i}\,(k_{y}\,y+k_{z}\,z)}\,\mathrm{d}k_{y}\,\mathrm{d}k_{z},$ (37)
where a truncation value of $\pm 5/\delta$ is used for the two wavenumbers in the computed integrals since the initial wavenumber $k_{0}=\sqrt{3}/\delta$ (Majumder et al., 2024). The evolution of blobs into columnar structures through the propagation of damped waves is evident in the yzy-z plane at $x=0$.
Figure 3: Variation of the squares of the fundamental frequencies with $Ra_{\ell,H}/Ra_{\ell}$ under constant vertical buoyancy conditions. Panels (a,c) and (b,d) correspond to cases with $\omega_{A,V}^{2}/\omega_{M}^{2}=0.8$ and $\omega_{A,V}^{2}/\omega_{M}^{2}=0.3$, respectively. In panels (a) and (b), $\omega_{C}^{2}$, $\omega_{M}^{2}$, and $\omega_{A}^{2}$ are plotted while panels (c) and (d) decompose $\omega_{A}^{2}$ into its vertical ($\omega_{A,V}^{2}$) and horizontal ($\omega_{A,H}^{2}$) parts. The dotted vertical lines indicate the values of $Ra_{\ell,H}/Ra_{\ell}$ where the slow MAC wave is suppressed as $|\omega_{M}|\approx|\omega_{A}|$. Panel (e) shows the suppression points of the slow wave for various strengths of $\omega_{A,V}^{2}/\omega_{M}^{2}$. The blue and red points in panel (e) correspond to the slow wave suppression points in panels (a) and (b), respectively. (f) The axial kinetic energy $E_{k,z}$ of the fast and slow MAC waves versus $Ra_{\ell,H}/Ra_{\ell}$ where $\omega_{A,V}^{2}/\omega_{M}^{2}=0.3$. The parameters used are $E_{\eta}=2\times 10^{-5}$, $Le=0.03$ ($\omega_{M}/\omega_{C}=0.18$), and $t/t_{\eta}=1\times 10^{-2}$.
In figure 3(a)–(d), the fundamental frequencies are plotted against the relative intensity of horizontal buoyancy. Both $Ra_{\ell,V}$ and $Le$ are kept constant. The frequencies and the local Rayleigh numbers are based on the mean wavenumbers obtained from ratios of $L^{2}$ norms; e.g.
[…]
By progressively increasing $Ra_{\ell,H}$ for a fixed $Ra_{\ell,V}$ and $Le$, $\omega_{A,H}$ increases, and the resultant buoyancy frequency $\omega_{A}$ crosses the Alfvén frequency $\omega_{M}$ at certain $Ra_{\ell,H}/Ra_{\ell}$, where the slow MAC waves of frequency $\omega_{s}$ disappear ($\omega_{s}=\omega_{3,4}$; see (29)). The same exercise performed with a different value of $Ra_{\ell,V}$ reveals the complementarity between $Ra_{\ell,V}/Ra_{\ell}$ and $Ra_{\ell,H}/Ra_{\ell}$ (figure 3(e)). The red and blue points in figure 3(e) indicate the value of $Ra_{\ell,H}/Ra_{\ell}$ from figure 3(a) and (b), where $|\omega_{A}|\approx|\omega_{M}|$ so that $\omega_{s}$ approaches zero. In §3, it is shown that the inequality $|\omega_{M}|>|\omega_{A}|$ represents the dipole-dominated regime while $|\omega_{M}|\approx|\omega_{A}|$ represents the transition to a multipolar state, with polarity reversals occurring in a narrow range of $Ra_{\ell,H}/Ra_{\ell}$ that lies between the dipolar and multipolar states.
The axial kinetic energy of the MAC waves is given by
$E_{k,z}=\frac{1}{(2\pi)^{2}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}|\hat{u}_{z}(k_{y},k_{z})|^{2}\mbox{d}k_{y}\mbox{d}k_{z},$ (39)
where the limits of the integration are set to $\pm 5/\delta$. While the energy of the fast MAC waves is practically unaffected by increasing horizontal buoyancy, the slow MAC wave energy goes to zero when $|\omega_{A}|$ approximately matches $|\omega_{M}|$ (figure 3(f)).
Figure 4: Variation of $Ra_{\ell}$ with Elsasser number $\varLambda=\left(\omega_{M}^{2}/\omega_{C}\omega_{\eta}\right)_{0}$ for the state of vanishing slow wave axial kinetic energy. Three values of $E_{\eta}$ are considered: diamonds represent $E_{\eta}=1\times 10^{-6}$, circles represent $E_{\eta}=2\times 10^{-5}$, and triangles represent $E_{\eta}=5\times 10^{-6}$.
The relation between the resultant local Rayleigh number $Ra_{\ell}$ and the ratio of frequencies given by
$\bigg(\dfrac{\omega_{M}^{2}}{\omega_{C}\omega_{\eta}}\bigg)_{\!0}\sim\dfrac{V_{M}^{2}}{2\varOmega\eta}$ (40)
is shown in figure 4 for the condition of vanishing slow wave axial kinetic energy for three different values of $E_{\eta}$. Each point is obtained by increasing the forcing through $Ra_{\ell}$ and noting the transition from the regime $|\omega_{M}|>|\omega_{A}|$ to the regime $|\omega_{A}|\gtrsim|\omega_{M}|$. The resultant forcing needed to suppress the slow MAC waves increases with $(\omega_{M}^{2}/\omega_{C}\omega_{\eta})_{0}$, which is a measure of the square of the peak dimensionless magnetic field in the dynamo (see §3.4 below). The linear relation obtained in figure 4 reflects the parity between $|\omega_{M}|$ and $|\omega_{A}|$ at the state of vanishing slow MAC waves. Since the relative orders of magnitude of the basic frequencies do not depend on the orientation of the gravity, rotation and magnetic field axes, the Cartesian model serves as the basis for the study of the role of equatorially anti-symmetric boundary heat flux in the nonlinear dynamo in §3, where the state of vanishing slow waves is shown to be a proxy for polarity transitions. We anticipate a self-similar behaviour for the dipole–multipole transition in the dynamo, where both $Ra_{\ell}$ and the square of the peak magnetic field are measurable quantities.
3 Nonlinear dynamo simulations
We consider an electrically conducting fluid confined between two concentric, co-rotating spherical surfaces. The ratio of the inner radius to the outer radius is 0.35. Fluid motion is driven by thermal buoyancy-driven convection. Lengths are scaled by the depth of the spherical shell $L$ and time is scaled by magnetic diffusion time $L^{2}/\eta$. The velocity $\bm{u}$ and magnetic field $\bm{B}$ are scaled by $\eta/L$ and $(2\Omega\rho\mu\eta)^{1/2}$ respectively. The temperature is scaled by $\beta_{s}L$, where $\beta_{s}$ is the mean equatorial radial temperature gradient within the shell. In the Boussinesq approximation, the nondimensional magnetohydrodynamic (MHD) equations for velocity, magnetic field, and temperature are as follows:
$EPm^{-1}\Bigl(\frac{\partial{\bm{u}}}{\partial t}+(\nabla\times{\bm{u}})\times{\bm{u}}\Bigr)+\hat{\bm{e}}_{z}\times{\bm{u}}=-\nabla p^{\star}+Ra_{V}\,PmPr^{-1}\,T\,{\bm{r}}\,$
$+(\nabla\times{\bm{B}})\times{\bm{B}}+E\nabla^{2}{\bm{u}},$ (41)
$\frac{\partial{\bm{B}}}{\partial t}=\nabla\times({\bm{u}}\times{\bm{B}})+\nabla^{2}{\bm{B}},$ (42)
$\frac{\partial T}{\partial t}+({\bm{u}}\cdot\nabla)T=PmPr^{-1}\,\nabla^{2}T,$ (43)
$\nabla\cdot{\bm{u}}=\nabla\cdot{\bm{B}}=0,$ (44)
The modified pressure $p^{*}$ in equation (41) is given by $p+\frac{1}{2}E\,Pm^{-1}\,|\bm{u}|^{2}$. The dimensionless parameters in the above equations are the Ekman number $E=\nu/2\varOmega L^{2}$, the Prandtl number, $Pr=\nu/\kappa$, the magnetic Prandtl number, $Pm=\nu/\eta$ and the modified vertical Rayleigh number $Ra_{V}=g\alpha\beta_{s}L^{2}/2\Omega\kappa$. Here, $g$ is the gravitational acceleration, $\nu$ is the kinematic viscosity, $\kappa$ is the thermal diffusivity and $\alpha$ is the coefficient of thermal expansion. The basic state conductive temperature profile is one of basal heating, given by $T_{0}(r)=r_{i}r_{o}/r$, where $r_{i}$ and $r_{o}$ are the inner and outer radii of the spherical shell. The velocity and magnetic fields satisfy the no-slip and electrically insulating conditions respectively at the two boundaries. The inner boundary is isothermal while a fixed heat flux at the outer boundary is given as the sum of the uniform mean and a lateral variation. The heterogeneity at the outer boundary is measured by the ratio
$q^{*}=\frac{\text{Maximum heat flux variation}}{\text{Mean heat flux}}.$ (45)
The mean basic state heat flux in the denominator of (45) represents the superadiabatic heat flux from the core (e.g. Olson, 2015, p. 11). The maximum heat flux variation at the outer boundary represents the magnitude of the peak-to-peak variation at the CMB relative to the mean superadiabatic value. The calculations are performed by a pseudospectral code that uses spherical harmonic expansions in the angular coordinates $(\theta,\phi)$ and finite differences in radius $r$ (Willis et al., 2007).
The parameter space in focus is that of low inertia, wherein the Rossby number based on the characteristic length scale of convection, $Ro_{\ell}$ (Christensen and Aubert, 2006) is small. Two Ekman numbers are considered, with the value of $Pm=Pr$ set to keep $Ro_{\ell}\ll 0.1$, ensuring that the simulations remain in the low-inertia regime. We explore the effect of inhomogeneous heat flux boundary conditions at the outer boundary on the dipole–multipole transition. To this end, we progressively increase the magnitude of the heat flux heterogeneity at the outer boundary for a given $Ra_{V}$. For a given $q^{*}$, the critical vertical Rayleigh number for the onset of nonmagnetic convection, $Ra_{V,\,c}$, is obtained for neutral stability of the perturbations subject to a steady mean flow and temperature field produced by the heterogeneity (Sahoo and Sreenivasan, 2017). The ratio $Ra_{V}/Ra_{V,\,c}$ (table 1) is a measure of the supercriticality of convection.
The mean spherical harmonic degrees for convection and energy injection, denoted by $l_{c}$ (Christensen and Aubert, 2006) and $l_{E}$ (Sreenivasan et al., 2014; Varma and Sreenivasan, 2022; Majumder et al., 2024) respectively, are given by,
$l_{c}=\dfrac{\Sigma\hskip 1.0ptl\hskip 1.0ptE_{k}(l)}{\Sigma\hskip 1.0ptE_{k}(l)};\hskip 5.0ptl_{E}=\dfrac{\Sigma\hskip 1.0ptl\hskip 1.0ptE_{T}(l)}{\Sigma\hskip 1.0ptE_{T}(l)},$ (46)
where $E_{k}(l)$ is the kinetic energy spectrum and $E_{T}(l)$ is the spectrum obtained from the product of the transform of $u_{r}T$ and its conjugate, showing the spectral distribution of scales at which energy is injected by buoyancy. The total kinetic and magnetic energies in the saturated dynamo are given by the volume integrals
$E_{k}=\dfrac{1}{2}\int\bm{u}^{2}\mbox{d}V;\quad E_{m}=\dfrac{Pm}{2E}\int\bm{B}^{2}\mbox{d}V.$ (47)
The relative dipole field strength $f_{dip}$, which is the ratio of the mean dipole field strength to the field strength in harmonic degrees $l=$ 1–12 at the outer boundary (Christensen and Aubert, 2006), takes values >0.35>0.35 in all the dipole-dominated runs (see table 1).
The square of the peak magnetic field, denoted by $B^{2}_{peak}$ (tables 1 & 2), is obtained from the time-averaged value of the magnetic field at the peak-field location in the saturated state of each run.
3.1 The effect of equatorially symmetric and anti-symmetric heat flux patterns on polarity transitions
(a)                (b)
Refer to caption Refer to caption
(c)                (d)
Refer to caption Refer to caption
Figure 5: The distribution of heterogeneous radial temperature gradient $\partial T/\partial r$ at the outer boundary for different conditions. (a) $Y_{2}^{1}$, (b) $Y^{2}_{2}$, (c) $Y_{2}^{2}:Y_{2}^{1}=2:1$ and (d) tomographic condition derived from the seismic shear wave velocity variation in the Earth’s lower mantle (Masters et al., 1996).
(a)                  (b)
Refer to caption Refer to caption
(c)                  (d)
Refer to caption Refer to caption
Figure 6: Evolution of dipole colatitude with magnetic diffusion time for (a) $q^{*}=17$, $Y_{2}^{1}$ (stable dipolar), (b) $q^{*}=18$, $Y_{2}^{1}$ (reversing), (c) $q^{*}=20$, $Y_{2}^{1}$ (multipolar) and (d) $q^{*}=18$ and $30$, $Y_{2}^{2}$ (stable dipolar). The other dynamo parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
(a)              (b)             (c)
Refer to caption Refer to caption Refer to caption
(d)              (e)             (f)
Refer to caption Refer to caption Refer to caption
Figure 7: Contours of the radial magnetic field at the outer boundary for $Ra_{V}=2500$ at $q^{*}=17$ (a), 18 (b), 20 (c); and for $Ra_{V}=25000$ at $q^{*}=5$ (d), 6 (e), 7 (f). The other dynamo parameters are $E=1.2\times 10^{-5}$, $Pm=Pr=1$. A $Y_{2}^{1}$ heat flux heterogeneity is applied at the outer boundary.
We begin our simulations with a strong field dynamo ($B^{2}_{rms}>1$) and increase the heterogeneity of the heat flux at the outer boundary in steps while keeping all other parameters fixed. Figure 5 (a)–(d) gives examples of the equatorially anti-symmetric, symmetric and composite boundary conditions whose effects on polarity transitions are analyzed in this study. The simulations are performed for at least one magnetic diffusion time in the saturated state.
In figure 6, the colatitude of the axial dipole field, $\theta$, at the outer boundary is obtained from the Gauss coefficients of the spherical harmonics, as follows:
$\cos\theta=g_{1}^{0}/|\bm{m}|,\quad\bm{m}=(g_{1}^{0},g_{1}^{1},h_{1}^{1}),$ (48)
where $g_{1}^{0}$, $g_{1}^{1}$, and $h_{1}^{1}$ are derived from the Schmidt-normalized expansion for the scalar potential of the field (Glatzmaier, 2013, pp. 142–143).
For $E=1.2\times 10^{-5}$, $Pm=Pr=1$, and $Ra_{V}=2500$, a strong dipole-dominated dynamo undergoes well-defined polarity reversals as shown in figure 6(b) by progressively increasing $q^{*}$ for the equatorially anti-symmetric $Y_{2}^{1}$ heat flux pattern. With a further increase in $q^{*}$, the dynamo exhibits a multipolar solution (figure 6(c)). However, the equatorially symmetric $Y_{2}^{2}$ heat flux pattern does not produce polarity transitions even with a strong heterogeneity (see figure 6(d)). For the same parameters, the Earth-like composite heat flux pattern in figure 5(d), which consists of both symmetric and anti-symmetric parts, produces polarity transitions. From the evolution of the dipole axis tilt (figure 17, Appendix A), well-defined polarity reversals are noted for $q^{*}=13$. For the anti-symmetric $Y_{2}^{1}$ boundary heat flux variation, the radial magnetic fields in the dipolar, reversing and multipolar states are shown in figure 7 for two series with different vertical thermal forcing. In table 1, the runs for two values of $Ra_{V}$, $2500$ and $20000$, are presented with different heat flux patterns. In the reversing runs, the ratio of magnetic to kinetic energy $E_{m}/E_{k}$ falls below unity, as noted in earlier studies (Kutzner and Christensen, 2002; Tassin et al., 2021) although a recent study (Frasson et al., 2025) suggests that multipolar solutions can exist for $E_{m}/E_{k}>1$. Reversing dynamo models in general have not realized $E_{m}/E_{k}>1$ probably because their magnetic Ekman number $E_{\eta}$ is much higher than that in the Earth’s core in the relatively small length scales. In contrast, the criterion of vanishing slow MAC wave helicity considered in this study applies to the energy-containing scales where density perturbations are continually produced, and is independent of the choice of $E_{\eta}$ in the low-inertia limit that rapidly rotating planets operate in.
The main output parameters of the polarity-reversing dynamo runs with the anti-symmetric $Y_{2}^{1}$ heat flux boundary condition are reported in table 2.
$q^{*}$ $\dfrac{Ra_{V}}{Ra_{V,\,c}}$ $N_{r}$ $l_{max}$ RmRm $Ro_{\ell}$ $l_{C}$ $l_{E}$ $\overline{m}$ $\bar{k}_{s}$ $\bar{k}_{z}$ $E_{k}$ $E_{m}$ $f_{dip}$ $B^{2}_{peak}$ $B^{2}_{rms}$ $\bigg|\dfrac{\omega_{A}}{\omega_{M}}\bigg|$ $\dfrac{Ra_{\ell,H}}{Ra_{\ell}}$ Type
$\times 10^{5}$ $\times 10^{5}$ D/R/M
$Ra_{V}=2500$, $Y_{2}^{1}$
0 75.75 168 160 174 0.011 19 26 7 4.04 2.82 2.22 27.25 0.81 95 3.59 0.49 0.00 D
5 80.64 176 176 181 0.013 19 28 7 4.46 4.43 2.43 26.30 0.77 81 3.51 0.60 0.36 D
10 80.64 176 176 232 0.018 19 28 7 4.41 3.94 4.06 16.80 0.71 78 2.26 0.76 0.51 D
15 80.64 176 176 250 0.020 19 28 7 4.01 3.92 4.81 14.16 0.67 72 1.66 0.86 0.56 D
17 80.64 176 176 261 0.021 19 28 7 4.39 3.97 5.15 13.11 0.65 67 1.52 0.92 0.60 D
18 80.64 176 176 341 0.025 19 28 7 4.55 3.92 8.58 2.41 0.25 41 0.40 1.00 0.62 R
20 80.64 176 176 349 0.027 19 27 7 3.93 3.83 8.93 2.43 0.13 43 0.38 1.19 0.64 M
$Ra_{V}=2500$, $Y_{2}^{2}$
18 125.00 176 176 239 0.018 20 34 7 4.45 4.90 4.15 15.68 0.82 75 2.56 0.42 D
25 131.57 176 176 251 0.017 18 35 6 4.74 4.99 4.60 17.68 0.81 88 2.82 0.41 D
30 138.88 176 176 263 0.018 18 35 5 4.59 5.02 5.03 18.68 0.78 90 3.01 0.41 D
$Ra_{V}=2500$, $Y_{2}^{2}:Y_{2}^{1}=2:1$
5 78.12 176 176 199 0.016 21 29 8 4.67 3.31 2.93 17.81 0.78 72 2.92 0.53 0.31 D
7 78.12 132 144 240 0.023 22 31 8 4.86 3.44 4.20 7.86 0.79 58 1.45 0.60 0.38 D
9 78.12 176 176 280 0.026 22 32 8 4.98 3.53 6.54 1.32 0.40 46 0.36 0.76 0.43 D
10 78.12 176 176 301 0.027 22 33 8 5.22 3.43 6.62 1.25 0.33 39 0.20 1.01 0.50 R
12 78.12 176 176 328 0.029 22 33 8 5.17 3.41 7.93 0.68 0.31 27 0.11 1.23 0.51 R
$Ra_{V}=2500$, $Y_{2}^{2}:Y_{2}^{1}=5:1$
10 80.64 132 144 222 0.017 21 34 7 3.43 3.83 3.47 18.83 0.69 86 2.65 0.30 0.24 D
30 80.64 176 176 264 0.018 22 34 7 3.82 4.11 5.07 15.01 0.79 116 2.51 0.41 0.44 D
$Ra=2500$, variation derived from seismic tomography
5 113.63 176 176 182 0.012 17 27 6 5.27 4.36 2.43 24.13 0.64 64 3.86 0.54 0.19 D
10 113.63 132 144 199 0.012 16 30 6 5.12 4.25 2.88 21.30 0.52 70 3.46 0.62 0.33 D
12 113.63 132 144 217 0.013 16 31 6 5.3 3.41 3.49 17.50 0.44 58 2.92 0.82 0.43 D
13 113.63 132 144 270 0.025 25 41 6 5.5 3.51 6.62 0.12 0.32 15 0.23 1.03 0.50 R
$Ra_{V}=20000$, $Y_{2}^{1}$
0 606.06 192 180 573 0.052 25 35 10 4.43 3.12 23.84 33.57 0.78 253 5.46 0.64 D
5 645.16 176 176 592 0.052 21 30 8 4.25 4.03 27.62 32.81 0.65 243 5.43 0.85 0.38 D
10 645.16 176 176 622 0.051 21 30 8 4.23 4.08 28.46 32.58 0.57 235 5.35 0.94 0.49 D
11 645.16 176 176 687 0.054 19 28 7 4.31 4.01 34.51 25.53 0.46 231 3.94 0.98 0.52 D
12 645.16 176 176 785 0.054 17 27 6 5.25 4.33 42.48 20.32 0.16 199 3.49 1.10 0.55 R
13 645.16 176 176 773 0.054 18 29 9 4.56 4.00 43.61 20.04 0.08 207 2.99 1.18 0.57 M
$Ra_{V}=20000$, $Y_{3}^{2}$
4 645.16 176 176 607 0.053 23 31 9 4.14 4.09 27.21 30.32 0.77 265 4.88 0.92 0.22 D
5 645.16 176 176 685 0.058 22 31 9 4.00 4.03 34.53 16.98 0.18 208 2.79 1.09 0.25 R
$Ra_{V}=20000$, $Y_{1}^{1}$
30 869.56 176 176 810 0.061 19 44 5 5.85 4.07 47.80 18.75 0.52 154 2.93 0.85 D
$Ra_{V}=20000$, $Y_{2}^{2}$
12 952.38 176 176 696 0.060 25 39 8 4.62 4.41 34.61 12.95 0.82 142 2.42 0.83 D
30 1111.11 176 176 776 0.073 24 40 7 4.61 4.76 44.39 7.80 0.57 104 1.26 0.92 D
$Ra_{V}=20000$, $Y_{3}^{3}$
30 952.38 176 176 782 0.072 25 37 8 4.23 4.45 37.62 8.77 0.52 94 0.98 0.95 D
$Ra_{V}=20000$, $Y_{3}^{1}$
30 645.16 176 176 782 0.07 25 38 8 4.18 4.52 31.44 35.61 0.54 156 3.18 0.77 D
Table 1: Summary of the main input and output parameters of a few dynamo simulations. Here, $Ra_{V}$ is the modified vertical Rayleigh number, $Ra_{V,\,c}$ is the modified vertical critical Rayleigh number for onset of nonmagnetic convection, $q^{*}$ is a dimensionless measure of the boundary heterogeneity defined in (45), $N_{r}$ is the number of radial grid points, $l_{\max}$ is the maximum spherical harmonic degree, RmRm is the magnetic Reynolds number, $Ro_{\ell}$ is the local Rossby number, $l_{C}$ and $l_{E}$ are the mean spherical harmonic degrees of convection and energy injection respectively, $\overline{m}$ is the mean spherical harmonic order in the range $l\leq l_{E}$, $\bar{k}_{s}$ and $\bar{k}_{z}$ are the mean $s$ and $z$ wavenumbers in the range $l\leq l_{E}$, $E_{k}$ and $E_{m}$ are the time-averaged total kinetic and magnetic energies, $f_{dip}$ is the relative dipole field strength, $Ra_{\ell,H}/Ra_{\ell}$ is the ratio of the local horizontal to the resultant Rayleigh numbers (obtained from (55) and (56)), and $B^{2}_{rms}$ is the measured mean square value of the field in the spherical shell. Types ‘D’, ‘R’, and ‘M’ denote dipolar, reversing and multipolar dynamo states, respectively. The other dynamo parameters are $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
$Ra_{V}$ $\dfrac{Ra_{V}}{Ra_{V,\,c}}$ $q^{*}$ $N_{r}$ $l_{max}$ RmRm $Ro_{\ell}$ $l_{C}$ $l_{E}$ $\overline{m}$ $\bar{k}_{s}$ $\bar{k}_{z}$ $E_{k}$ $E_{m}$ $f_{dip}$ $B^{2}_{rms}$ $Ra_{\ell}$ $B^{2}_{peak}$ $\dfrac{Ra_{\ell,H}}{Ra_{\ell}}$
$\times 10^{5}$ $\times 10^{5}$ $\times 10^{4}$
$E=6\times 10^{-5},Pm=Pr=5$
2000 54.05 25 176 176 235 0.012 14 21 5.0 3.74 3.81 4.03 1.05 0.11 0.12 0.71 85 0.66
3000 81.08 25 176 176 235 0.012 14 21 5.2 3.65 3.93 6.39 3.51 0.12 0.24 0.97 130 0.65
4500 121.62 22 176 176 333 0.018 14 21 5.3 3.6 3.61 8.04 2.07 0.08 0.35 1.25 160 0.64
6000 162.16 20 176 176 378 0.020 14 21 5.4 3.87 3.83 10.50 2.18 0.08 0.27 1.63 170 0.63
8000 216.21 18 176 176 427 0.023 14 21 5.6 3.89 3.98 13.17 2.35 0.07 0.38 2.01 210 0.61
10000 270.27 18 176 176 479 0.026 14 21 5.6 3.94 4.36 15.55 4.58 0.08 0.44 2.48 225 0.60
12000 324.32 17 176 176 509 0.028 15 22 5.8 4.25 4.58 18.83 4.83 0.12 0.59 2.84 250 0.58
15000 405.40 17 176 176 562 0.030 14 21 5.4 4.2 4.75 23.02 5.54 0.11 0.90 3.55 274 0.57
18000 486.48 14 176 176 603 0.033 14 21 5.6 4.27 4.35 26.38 6.20 0.11 1.04 3.85 295 0.54
20000 540.54 5 176 176 629 0.034 14 21 6.0 4.32 3.94 26.64 8.26 0.25 1.29 3.22 268 0.41
21000 552.63 0 160 180 549 0.039 19 25 8.1 4.22 3.38 20.54 11.47 0.30 0.38 2.25 227 0.00
$E=1.2\times 10^{-5},Pm=Pr=1$
1500 48.38 18 132 128 275 0.020 19 27 5.5 4.73 3.95 5.49 1.38 0.23 0.25 0.42 43 0.64
2500 80.64 18 176 176 341 0.025 19 28 7.0 4.55 3.92 8.58 2.41 0.25 0.40 0.69 65 0.63
4000 129.03 17 176 176 423 0.030 19 28 7.1 4.49 3.98 12.82 3.94 0.22 0.67 1.04 91 0.61
6000 193.54 17 176 176 485 0.033 19 28 7.1 4.55 4.21 16.90 6.82 0.15 1.16 1.52 120 0.60
8000 258.06 17 176 176 546 0.037 18 28 6.7 4.66 4.58 21.77 8.51 0.14 1.21 2.00 137 0.58
10000 322.58 17 176 176 605 0.040 18 28 6.5 4.39 4.49 26.52 10.70 0.10 1.74 2.42 169 0.57
12000 387.09 16 176 176 639 0.044 18 28 6.5 4.53 4.12 29.82 13.17 0.09 2.01 2.77 186 0.57
15000 483.87 15 176 176 703 0.045 17 27 6.1 4.83 4.28 34.77 16.65 0.11 2.68 3.23 202 0.56
20000 645.16 12 176 176 785 0.054 17 27 6.2 5.25 4.33 42.48 20.32 0.16 3.49 4.06 221 0.53
25000 806.45 6 176 176 795 0.060 18 28 7.2 5.18 4.14 44.83 21.34 0.11 1.85 3.45 217 0.42
27000 870.96 1 176 176 842 0.062 19 29 7.8 5.10 3.94 39.66 21.51 0.08 3.11 3.11 200 0.20
28000 848.48 0 192 180 775 0.073 25 36 10.0 4.81 2.91 39.74 19.67 0.25 0.67 2.80 197 0.00
Table 2: Summary of the main input and output parameters in the reversing dynamo simulations considered in this study for $Y_{2}^{1}$ heat flux condition. These simulations have $|\omega_{A}/\omega_{M}|\geq 1$ throughout. Here, $Ra_{V}$ is the modified Rayleigh number, $Ra_{V,\,c}$ is the modified vertical critical Rayleigh number for onset of nonmagnetic convection, $q^{*}$ is a dimensionless measure of the boundary heterogeneity defined in (45), $N_{r}$ is the number of radial grid points, $l_{max}$ is the maximum spherical harmonic degree, RmRm is the magnetic Reynolds number, $Ro_{\ell}$ is the local Rossby number, $l_{C}$ and $l_{E}$ are the mean spherical harmonic degrees of convection and energy injection respectively (defined in (46)), $\overline{m}$ is the mean spherical harmonic order in the range $l\leq l_{E}$, $\bar{k}_{s}$ and $\bar{k}_{z}$ are the mean $s$ and $z$ wavenumbers in the range $l\leq l_{E}$, $E_{k}$ and $E_{m}$ are the time-averaged total kinetic and magnetic energies defined in (47), $f_{dip}$ is the relative dipole field strength, $Ra_{\ell}$ is the local Rayleigh number defined in (55) and $B^{2}_{peak}$ is the square of the measured peak field when slow MAC waves cease to exist when $|\omega_{A}|\approx|\omega_{M}|$.
The ratio of anti-symmetric to total kinetic energy remains nearly constant even at high $q^{*}$ (figure 8), so equatorially anti-symmetric boundary conditions do not induce polarity transitions by breaking the equatorial symmetry of the convection columns. Before analysing the role of wave motions in polarity transitions, we examine the mean temperature and velocity fields produced by the heterogeneous boundary heat flux, which give useful comparisons with the linear convection model.
Figure 8: Ratio of anti-symmetric to total kinetic energy for $q^{*}=0$ (black), $q^{*}=15$ (red), $q^{*}=18$ (blue) and $q^{*}=20$ (green) with the $Y_{2}^{1}$ heat flux heterogeneity at the outer boundary. The other dynamo parameters are $Ra=2500,E=1.2\times 10^{-5},Pm=Pr=1$.
3.2 Mean resultant temperature gradient and velocity field
In line with that obtained in equation (36) in the Cartesian model, the mean (superadiabatic) resultant temperature gradient is given by
$\beta=\beta_{s}-2\beta_{z}k_{s}/k_{z},$ (49)
where $\beta_{s}=\partial T_{0}/\partial s<0$ in an unstably stratified layer with homogeneous boundary heat flux (see figure 9(a)). In figure 9(b), $\beta_{z}=\partial T_{0}/\partial z$ is plotted for an equatorially antisymmetric $Y_{2}^{1}$ heat-flux heterogeneity at the outer boundary. The temperature gradients are calculated from a steady state just below nonmagnetic convective onset ($Ra_{V}=30$, whereas onset occurs at $Ra_{V}=31$ for $q^{*}=18$) so that the perturbations do not affect the calculation of the mean gradients. An averaged $\beta_{z}$, obtained by taking the average of the positively signed $\partial T_{0}/\partial z$ within the spherical shell, is proportional to the intensity of the anti-symmetric heterogeneity. Likewise, the averaged $\beta_{s}$ is obtained by taking the average of $\partial T_{0}/\partial s$ within the shell. The averaged resultant gradient $\beta$ is then evaluated from (49) using the average $s$ and $z$ wavenumbers defined in (53) (see figure 9(d)). The resultant $\beta$ calculated at $Ra_{V}=30$ and the resultant $\beta$ averaged over $\approx$ 0.5 diffusion time from the strongly supercritical state at $Ra_{V}=310$ exhibit similar structures as well as nearly equal maximum and minimum values (see figure S4 in the Supplementary Material). Convection is suppressed in stably stratified regions where $\partial T/\partial r>0$. In contrast, regions beneath $\partial T/\partial r<0$, where $\beta_{z}>0$ and $\beta<0$, support convection. This unstably stratified region is in focus in the analysis in §3.3 below – at moderately positive values of $\beta_{z}$, both fast and slow MAC waves are present, giving the dipolar dynamo regime; at large positive values of $\beta_{z}$, the slow MAC waves are selectively suppressed whereas the fast MAC waves persist, giving the reversing and multipolar dynamo regimes.
Equation (49) is derived from the theoretical model described in §2, which predicts the presence or absence of slow MAC waves subject to the frequency inequality $|\omega_{C}|>|\omega_{M}|,|\omega_{A}|>|\omega_{\eta}|$, with slow MAC waves present when $|\omega_{M}|>|\omega_{A}|$ and absent when $|\omega_{M}|\leq|\omega_{A}|$. Since the above frequency inequality does not predict the localized suppression of convection observed in figure 9(f), equation (49) can only be applied to regions where $\beta_{z}>0$. Interestingly, $\beta_{\phi}=\partial T_{0}/\partial\phi$ (figure 9(c)) has practically no influence on $\beta$, which justifies our original assumption of ignoring variations with respect to the non-preferred horizontal direction (see (15), §2.2). The convective region produced by $\beta_{s}<0$ and $\beta_{z}>0$ with the equatorially anti-symmetric heat flux at the boundary corresponds to the unstably stratified system modelled by the equatorial radial configuration of the Cartesian linear model in §2.1.
Figure 10(a) shows the azimuthal variation of the mean temperature gradients at cylindrical radius $s=1$ and at the section $z=0.4$ below the equator for the $Y_{2}^{1}$ heterogeneity with $q^{*}=18$. The grey-shaded region indicates $\beta_{z}>0$, where the resultant $\beta$ is calculated from (49). Figure 10(b) gives the meridional plot of the steady mean flow $u_{0}$ ($\phi$-component of the velocity) averaged over the unstably stratified longitudes in figure 10(a) ($\phi=1.31\,\pi$ to $0.31\,\pi$ through $\phi=0$). In figure 10(c), the magnitude of this equatorially anti-symmetric flow measured at $s=1$ shows a fair agreement with the theoretical value of the thermal wind calculated from equation (4) in dimensionless units,
$u_{0}=-\frac{g\alpha\beta_{z}L}{2\varOmega\eta}\,z=-Ra_{V}\,\beta^{*}z,$ (50)
where $\beta^{*}=\beta_{z}/\beta_{s}$ is obtained from the respective averaged temperature gradients.
(a)                  (b)
Refer to caption Refer to caption
(c)                  (d)
Refer to caption Refer to caption
(e)                  (f)
Refer to caption Refer to caption
Figure 9: Horizontal section plots at $z=0.4$ below the equator showing (a) $\beta_{s}$ for homogeneous boundary heating, (b) $\beta_{z}$ for $q^{*}=18$, (c) $\beta_{\phi}$ for $q^{*}=18$, and (d) resultant gradient $\beta$ for $q^{*}=18$. The temperature gradients are calculated at vertical Rayleigh number $Ra_{V}=30$, which is a state near the onset of nonmagnetic convection. A $Y_{2}^{1}$ heterogeneity in outer boundary heat flux is applied. The orange-green coloured strip at the periphery of plot (b) represents $\partial T_{0}/\partial r$ at the outer boundary. Panels (e) and (f) show snapshots of the axial velocity $u_{z}$ at $q^{*}=0$ and $q^{*}=18$ respectively for dynamo simulations at $Ra_{V}=2500$. The other parameters are $E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
(a)              (b)         (c)
Refer to caption Refer to caption Refer to caption
Figure 10: (a) Values of $\beta_{s}$ (blue), $\beta_{z}$ (red) and the resultant $\beta$ (black) at cylindrical radius $s=1$ at the section $z=0.4$ below the equator for $q^{*}=18$. (b) Meridional sections plot of the $\phi$-component of the velocity, $u_{0}$ for $q^{*}=18$. The plot shows $\phi$-averaged values of $u_{0}$ in the unstably stratified region where $\beta<0$ ($\phi=1.31\,\pi$ to $0.31\,\pi$ through $\phi=0$) in the grey-shaded region of figure (a). (c) The values of $u_{0}$ along a vertical line passing through cylindrical radius $s=1$ (circles) are compared with the theoretical dimensionless thermal wind in equation (50). The parameters are $Ra_{V}=30$ (below onset), $E=1.2\times 10^{-5}$, and $Pm=Pr=1$. A $Y_{2}^{1}$ heat flux heterogeneity is applied at the outer boundary.
3.3 The role of slow MAC waves in the dipole–multipole transition
In §2, we examined the evolution of an isolated density perturbation in an unstably stratified fluid subject to background rotation, a uniform magnetic field and a horizontal temperature gradient. Perturbations of this kind excite MAC waves in the dynamo, the frequencies of which depend on the fundamental frequencies given below:
$\omega_{C}^{2}=\frac{4(\bm{\varOmega}\cdot\bm{k})^{2}}{k^{2}},\quad\omega_{M}^{2}=\frac{(\bm{B}\cdot\bm{k})^{2}}{\mu\rho},\quad-\omega_{A,V}^{2}=g\alpha\beta_{s}\,\frac{k_{h}^{2}}{k^{2}},\quad\omega_{A,H}^{2}=g\alpha\beta_{z}\frac{k_{h}^{2}}{k^{2}}\frac{k_{s}}{k_{z}}$ (51)
and scaling the frequencies by $\eta/L^{2}$, we obtain in dimensionless units,
$\omega_{C}^{2}=\frac{Pm^{2}}{E^{2}}\frac{k_{z}^{2}}{k^{2}},\quad\omega_{M}^{2}=\frac{Pm}{E}(\bm{B}\cdot{\bm{k}})^{2},\quad-\omega_{A,V}^{2}=\frac{PmRa_{V}}{E}\,\frac{{k_{h}}^{2}}{k^{2}},\quad\omega_{A,H}^{2}=\frac{PmRa_{H}}{2\,E}\,\frac{{k_{h}}^{2}}{k^{2}},\quad$ (52)
where $k_{s}$, $k_{\phi}$ and $k_{z}$ are the radial, azimuthal and axial wavenumbers in cylindrical coordinates $(s,\phi,z)$, $k_{\phi}=m/s$, where $m$ is the spherical harmonic order, $k^{2}=k_{s}^{2}+k_{\phi}^{2}+k_{z}^{2}$, and $k_{h}$ is the horizontal wavenumber in the equatorial region defined by $k_{h}^{2}=k_{\phi}^{2}+k_{z}^{2}$. The $s$, $\phi$ and $z$ wavenumbers are calculated in the saturated state of the dynamo runs in the energy-containing scales ($l\leq l_{E}$). For example, real space integration over $(s,\phi)$ gives the kinetic energy as a function of $z$, the Fourier transform of which gives the one-dimensional spectrum $\hat{u}^{2}(k_{z})$. Subsequently, we obtain,
$\bar{k}_{z}=\frac{\Sigma k_{z}\,\hat{u}^{2}(k_{z})}{\Sigma\hat{u}^{2}(k_{z})}.$ (53)
A similar approach yields $\bar{k}_{s}$ and $\overline{m}$. The magnetic (Alfvén) wave frequency $\omega_{M}$ is based on the three components of the measured magnetic field at the peak-field location. The wavenumber $k_{\phi}$ is evaluated at $s=1$, approximately mid-radius of the spherical shell. The horizontal Rayleigh number is given by
$Ra_{H}=2\,\frac{g\alpha\beta_{z}L^{2}}{2\Omega\eta}\frac{k_{s}}{k_{z}}.$ (54)
The local vertical and horizontal Rayleigh numbers in the dynamo based on the flow length scale are given by,
$Ra_{\ell,V}=\frac{g\alpha|\beta_{s}|}{2\Omega\eta}\left(\frac{\pi}{k_{s}}\right)^{2},\,Ra_{\ell,H}=2\,\frac{g\alpha|\beta_{z}|}{2\Omega\eta}\frac{k_{s}}{k_{z}}\left(\frac{\pi}{k_{s}}\right)^{2},$ (55)
and the resultant local Rayleigh number is then
$Ra_{\ell}=Ra_{\ell,V}+Ra_{\ell,H}.$ (56)
Figure 11: Variation of the squares of the fundamental frequencies with $q^{*}$ and $Ra_{\ell,H}/Ra_{\ell}$ (within brackets) for (a) $Ra_{V}=2500$, $Y_{2}^{1}$ heat flux heterogeneity, and (b) $Ra_{V}=2500$, $Y_{2}^{2}$ heat flux heterogeneity. The dotted vertical line marks the polarity-reversing state that lies between the dipolar and multipolar regimes. The other dynamo parameters are $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
Figure 12: Variation of $f_{dip}$ with $|\omega_{A}/\omega_{M}|$ for different heat flux boundary conditions at two values of $Ra_{V}$. The vertical dotted line indicates $|\omega_{A}/\omega_{M}|=1$, while the horizontal dotted line marks $f_{dip}=0.35$, which has been proposed as the lower bound for the existence of dipole-dominated numerical dynamos (Christensen and Aubert, 2006). Types ‘D’, ‘R’, and ‘M’ denote dipolar, reversing and multipolar dynamo states, respectively. The other dynamo parameters are $E=1.2\times 10^{-5}$ and $Pm=Pr=1$.
In figure 11, the magnitudes of fundamental frequencies are shown as a function of $q^{*}$ and $Ra_{\ell,H}/Ra_{\ell}$ for heterogeneous heat flux boundary conditions. The frequencies and local Rayleigh numbers are computed from (52), (55) and (56) using the mean values of the wavenumbers as shown in equation (53). In Figure 11(a), we begin with a strong field dynamo and increase the $Y_{2}^{1}$ heat flux heterogeneity, which is reflected in the horizontal buoyancy frequency $\omega_{A,H}$, while keeping the vertical buoyancy constant at $Ra_{V}=2500$. The slow MAC waves are present at relatively low $q^{*}$ when $|\omega_{M}|>|\omega_{A}|$. In the polarity-reversing state at $q^{*}=18$ ($Ra_{\ell,{H}}/Ra_{\ell}\approx 0.62$), the magnitude of the resultant buoyancy frequency $|\omega_{A}|$ approximately matches that of the Alfvén frequency $|\omega_{M}|$, causing the slow wave frequency $\omega_{s}$ to go to zero. While the generation of slow MAC waves ceases, the existing slow MAC waves are damped on the time scale $\delta^{2}/\eta$ (see §2.4), which is much shorter than the magnetic diffusion time $L^{2}/\eta$ because the ratio of the core depth to the length scale of buoyancy disturbances would be $L/\delta\sim 10^{2}$ (Sreenivasan and Maurya, 2021). Whether the rapid damping of the slow MAC waves causes the rapid decay of the axial dipole field during a polarity reversal (see figure 6 b) is at present an open question that requires further investigation. The equatorially symmetric $Y_{2}^{2}$ heat flux boundary condition has zero horizontal buoyancy at the equator, and even at large $q^{*}$, the condition $|\omega_{A}|>|\omega_{M}|$ is never met (figure 11(b)).
The variation of the parameter $f_{dip}$ against $|\omega_{A}/\omega_{M}|$ in figure 12 shows that reversing (R) and multipolar (M) states exist at values of $f_{dip}\leq 0.35$, the proposed lower bound for the existence of dipole-dominated numerical dynamos in the literature.
Figure 13: Panels (a)–(c) (i) show the absolute values of wave frequencies plotted for the saturated state of the dynamo run. (a) $q^{*}=0$, (b) $q^{*}=10$ and (c) $q^{*}=18$ for the $Y_{2}^{1}$ heat flux heterogeneity at the outer boundary. The shaded grey area shows the range of scales where the helicity of the dynamo run is greater than that of the equivalent non-magnetic run. Panels (a)–(c) (ii) show the spectral distribution of the power supplied to the axial dipole, defined in (57). The other dynamo parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
The top panel of figure 13 displays the square of the fundamental frequencies for a range of spherical harmonic order $m$ for (a) $q^{*}=0$, (b) $q^{*}=10$, and (c) $q^{*}=18$ in the saturated state for the equatorially anti-symmetric $Y_{2}^{1}$ heat flux pattern. Dynamo simulations with $q^{*}=0$ and $q^{*}=10$ exhibit stable axial dipoles while the simulation with $q^{*}=18$ gives a reversing dynamo. The shaded area indicates the range of $m$ where the kinetic helicity in the dynamo exceeds that of the equivalent non-magnetic run. (For the governing equations of equivalent non-magnetic run, see Majumder et al., 2024). The black line represents the square of the slow MAC wave frequency under the condition $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|$. As $q^{*}$ increases, the range of $m$ satisfying this condition narrows, and for $q^{*}=18$, this condition is not satisfied for any $m$. The bottom panel of figure 13 gives the spectral distribution of the power supplied to the poloidal component of the axial dipole field $B_{10}^{P}$ (e.g. Buffett and Bloxham, 2002) for the three $q^{*}$ values,
$P_{\mathrm{10}}=\int_{V}\bm{B}^{P}_{\mathrm{10}}\cdot[\mathbf{\nabla}\times(\bm{u}\times\bm{B})_{m}]\,dV,$ (57)
where $\bm{u}$ and $\bm{B}$ have the same $m$ (Bullard and Gellman, 1954). The dipole power peaks at $q^{*}=0$ and $q^{*}=10$ within the range of $m$ where slow MAC waves are generated and the helicity of the dynamo is greater than that of the equivalent non-magnetic simulation. In contrast, the run with $q^{*}=18$ shows no such peak since there is no helicity generation at any $m$. The $f_{dip}$ values for $Ra=2500$ decrease with increasing heterogeneity, with a sharp drop observed at the polarity transition for $q^{*}=18$ (table 1). These results indicate the crucial role of the slow MAC waves in helicity generation in the energy-containing scales, and in turn, dipole formation.
(a)                  (b)
Refer to caption Refer to caption
Figure 14: (a-b) Contour plots of $\partial{u}_{z}/\partial t$ at a cylindrical radius of $s=1$ are shown for the scales $l\leq l_{E}$ over short time intervals in the saturated state of two dynamo simulations with (a) $q^{*}=17$ and (b) $q^{*}=18$ for the equatorially anti-symmetric $Y_{2}^{1}$ heat flux boundary condition. The parallel black lines represent the primary wave travel direction, with their slope giving the group velocity $U_{g,z}$. Table 3 lists the estimated group velocities for the fast and slow MAC waves ($U_{f}$ and $U_{s}$, respectively) and $U_{g,z}$. The other dynamo parameters are $Ra_{V}=2500,E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
SL $E$ $Ra_{V}$ $q^{*}$ Fig. $\omega_{n}^{2}$ $\omega_{C}^{2}$ $\omega_{M}^{2}$ ${\omega_{A}}^{2}$ $\omega_{f}$ $\omega_{s}$ $U_{f}$ $U_{s}$ $U_{g,z}$
No. $(\times 10^{10})$ $(\times 10^{8})$ $(\times 10^{8})$ $(\times 10^{8})$ $(\times 10^{4})$ $(\times 10^{4})$
1 $1.2\times 10^{-5}$ 2500 17 14 (a) 2.5 13.02 6.59 3.94 4.63 0.90 $89$ 525 456
2 $1.2\times 10^{-5}$ 2500 18 14 (b) 2.7 11.51 2.22 4.32 3.46 0.0 9651 0 9274
Table 3: Summary of the data for MAC wave measurement in the dynamo models. The sampling frequency $\omega_{n}$ is selected to ensure that the fast MAC waves are captured when measuring group velocity. The values of $\omega_{M}^{2}$, $\omega_{A}^{2}$, and $\omega_{C}^{2}$ are computed using (52), based on the averaged wavenumbers $m$, $k_{s}$, and $k_{z}$ in the energy-containing scales where $l\leq l_{E}$. The group velocity in the $z$ direction ($U_{g,z}$) is then compared with the estimated velocities of the fast ($U_{f}$) and slow ($U_{s}$) MAC waves.
Figure 14 shows the measurement of wave velocities from their propagation paths in the saturated state of dynamos at $q^{*}=17$ and $q^{*}=18$ for the equatorially anti-symmetric $Y_{2}^{1}$ heat flux pattern at the outer boundary. Contours of the fluctuating $z$-velocity, represented by $\partial u_{z}/\partial t$ at cylindrical radius $s=1$, are plotted over short time intervals where the ambient magnetic field and wavenumbers in the energy-containing scales $l\leq l_{E}$ remain approximately constant. By examining the slope of the black lines, we determine the axial group velocity $U_{g,z}$ and compare it with the estimated axial group velocities of the fast ($U_{f}$) and slow ($U_{s}$) waves. These velocities are derived from the respective frequencies in the diffusionless limit by taking their derivative with respect to $k_{z}$ (Varma and Sreenivasan, 2022). In the dipole-dominated run at $q^{*}=17$, the slow MAC waves are predominant although the fast MAC waves also exist (figure 14(a)). In contrast, at $q^{*}=18$, the slow MAC waves are nearly absent while the fast waves are abundant (figure 14(b)).
3.4 Complementarity of vertical and horizontal buoyancies in the dipole–multipole transition
(a)                  (b)
Refer to caption Refer to caption
(c)                  (d)
Refer to caption Refer to caption
Figure 15: (a) Variation of the modified vertical Rayleigh number $Ra_{V}$ with the square of the peak magnetic field, $B^{2}_{peak}$ at the suppression of slow MAC waves. (b) Variation of the local Rayleigh number $Ra_{\ell}$, defined in (55), with $B^{2}_{peak}$. The values of $_{V}, RaRa_{\ell}$ and $B^{2}_{peak}$ in the plots are given in table 2. The parameters of the two dynamo series and their symbolic representations are as follows: $E=6\times 10^{-5},Pm=Pr=5$ (circles), $E=1.2\times 10^{-5},Pm=Pr=1$ (diamonds). The filled symbols show the evolution path of the dynamo with increasing heat flux heterogeneity. Black symbols represent the dipolar state and red symbols represent the reversing or multipolar state. (c) The states of polarity reversals are shown in a plot of local relative vertical versus relative horizontal Rayleigh numbers, normalized by the resultant local Rayleigh number $Ra_{\ell}$. For $E=1.2\times 10^{-5}$, reversals occur at $Ra_{V}=27000$ and $Ra_{\ell,H}/Ra_{\ell}=0.2$ (blue diamond) as well as at $Ra_{V}=1500$ and $Ra_{\ell,H}/Ra_{\ell}=0.64$ (red diamond). (d) Variation of $q^{*}$ with respect to $Ra_{\ell,H}/Ra_{\ell}$. The horizontal line at $q^{*}=10$ and the vertical line at $Ra_{\ell,H}/Ra_{\ell}=0.5$ indicate that $Ra_{\ell,H}/Ra_{\ell}>0.5$ corresponds to $q^{*}=O(10)$.
In the presence of rapid rotation and a magnetic field, the buoyant forcing generates fast and slow MAC waves when the inequality $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|>|\omega_{\eta}|$ is satisfied. The magnetic field intensity increases with the vigour of convection in this dipole-dominated regime; however, a sufficiently strong forcing suppresses the slow MAC waves when $|\omega_{A}|\approx|\omega_{M}|$, leading to dipole collapse. The nature of buoyancy could be either vertical, measured by $Ra_{V}$, or horizontal, measured by $Ra_{H}$. In a dynamo with homogeneous boundary heat flux, the dipolar regime transitions in succession to polarity-reversing and multipolar states with increasing $Ra_{V}$ (Majumder et al., 2024). In this study, $Ra_{H}$ is progressively increased for a fixed $Ra_{V}$ until the dynamo undergoes the polarity transition (table 1). The square of the peak magnetic field measured at the polarity transition, $B^{2}_{peak}$, is plotted against $Ra_{V}$ for two different Ekman numbers in figure 15(a). The approximately linear variation of $Ra_{V}$ with $B^{2}_{peak}$ is different in the two cases. On the other hand, the resultant local Rayleigh number $Ra_{\ell}$, defined by equation (56), exhibits a self-similar variation with $B^{2}_{peak}$ (figure 15(b)). This self-similar line separates the dipolar and multipolar regimes. The filled symbols represent the evolution path of the dynamo with increasing order of heat flux heterogeneity for two vertical Rayleigh numbers ($Ra_{V}=2500$ and $Ra_{V}=20000$). The filled black symbols here represent the states where $|\omega_{M}|>|\omega_{A}|$ while the filled red symbols represent $|\omega_{M}|\lesssim|\omega_{A}|$. In figure 15(c), all the polarity-reversing states are plotted in the $Ra_{\ell,H}$ versus $Ra_{\ell,V}$ space, normalized by the value of $Ra_{\ell}$. This plot also demarcates the dipolar and multipolar regimes and furthermore demonstrates the complementarity between the horizontal and vertical buoyancies in producing the polarity transition through the condition of vanishing slow MAC waves, as suggested by figure 3(e). For $E=1.2\times 10^{-5}$, reversals occur in a strongly driven dynamo at $Ra_{V}=27000$ and $Ra_{\ell,H}/Ra_{\ell}=0.2$ (blue diamond) as well as a relatively weakly driven dynamo at $Ra_{V}=1500$ and $Ra_{\ell,H}/Ra_{\ell}=0.64$ (red diamond); see also table 2. Since the self-generated magnetic field depends on the vertical Rayleigh number $Ra_{\ell,V}$, only weak-field dynamos of $B^{2}_{rms}<O(1)$ exist for $Ra_{\ell,V}/Ra_{\ell}<0.25$. Therefore, solutions for $Ra_{\ell,V}/Ra_{\ell}<0.25$ (i.e. $Ra_{\ell,H}/Ra_{\ell}>0.75$) are not considered. Figure 15(d) relates the two measures of heterogeneity, $q^{*}$ and $Ra_{\ell,H}/Ra_{\ell}$ and indicates that values of $Ra_{\ell,H}/Ra_{\ell}>0.5$ correspond to values of $q^{*}$ of $O(10)$.
3.5 Combination of symmetric and anti-symmetric boundary heat flux
Since the heat flux heterogeneity in the lower mantle is a combination of both equatorially symmetric and anti-symmetric variations, it is instructive to consider the effect of composite heat flux boundary conditions on polarity transitions. The analysis of the heat flux distribution based on the seismic shear wave velocity in the Earth’s lower mantle (Masters et al., 1996) suggests a complex heat flux pattern at the CMB which features a well-defined $Y_{2}^{2}$ component. That said, the analysis of peak-to-peak heat flux variations reveals that the ratio of the symmetric to anti-symmetric heat flux variation is approximately 1.65. This ratio is obtained by calculating the peak-to-peak variation in heat flux for the symmetric and anti-symmetric contributions separately, within the seismic tomography pattern. With a composite heat flux heterogeneity at the outer boundary, a dipole-dominated solution is observed when the inequality $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|$ is satisfied in the presence of the slow MAC waves. For a given vertical buoyant forcing ($Ra_{V}\sim 10^{2}$ times its value for nonmagnetic onset of convection), the progressive increase of horizontal buoyancy produces polarity transitions for composite patterns where the equatorially symmetric and anti-symmetric variations are comparable. For example, a heat flux pattern with $Y_{2}^{2}:Y_{2}^{1}=2:1$ and the Earth-like heat flux pattern derived from seismic tomography both produce reversals for $q^{*}$ of $O(10)$ (table 1). For the same $Ra_{V}$, the values of $Ra_{\ell,H}/Ra_{\ell}$ at which the transition occurs in the two cases are nearly equal, which points to the comparable ratios of the symmetric to anti-symmetric heat flux variation. Notably, a heterogeneity consisting of comparable magnitudes of symmetric and anti-symmetric variations induces the polarity transition at a value of $Ra_{\ell,H}/Ra_{\ell}$ (and $q^{*}$) which is of the same order as that for the transition induced by a purely anti-symmetric variation.
The polarity transition with the boundary heat flux derived from seismic tomography is analysed further in figure 18 of Appendix A. In the dipole-dominated state at $q^{*}=10$, $|\omega_{M}|^{2}>|\omega_{A}|^{2}$ in several regions, where $\omega_{A}^{2}$ is based on the resultant basic state temperature gradient $\beta$, defined in (49). In these regions, the Alfvén wave velocity is greater than the velocity of the buoyancy perturbations and the slow MAC waves are generated. On the other hand, the polarity-reversing state at $q^{*}=13$ shows approximate parity between $|\omega_{M}|^{2}$ and $|\omega_{A}|^{2}$ so that the slow waves are suppressed. This analysis also explains why a composite boundary heterogeneity made up of a dominant equatorially symmetric variation does not induce polarity transitions. For example, the heterogeneity with the ratio $Y_{2}^{2}:Y_{2}^{1}=5:1$ does not admit polarity transitions even at high $q^{*}$ (table 1) since the condition $|\omega_{A}|^{2}\gtrsim|\omega_{M}|^{2}$ is not met.
4 Two-component magnetoconvection
Convection in the cores of planets like Earth are driven by compositional and thermal buoyancy. The compositional part in Earth, arising from the progressive growth of the inner core, is dominant. With the knowledge of the peak magnetic field intensity, a useful constraint on the lateral variation in heat flux can be obtained from an analysis of the evolution of an isolated disturbance under rapid rotation and both compositional and thermal buoyancy.
4.1 Evolution of a density disturbance in two-component magnetoconvection
In addition to the variables $\bm{u}$, $\bm{b}$ and $\Theta$ considered in (7)(a–c), the composition $C$ is also decomposed into its mean and perturbation parts, $C=C_{0}+\gamma$. Here, the lengths are scaled by the perturbation size $\delta$ and time is scaled by magnetic diffusion time $\delta^{2}/\eta$. The velocity $\bm{u}$ and magnetic field $\bm{B}$ are scaled by $\eta/\delta$ and $(2\Omega\rho\mu\eta)^{1/2}$, respectively. The temperature is scaled by $\beta_{y}^{T}\delta$ and composition is scaled by $\beta_{y}^{C}\delta$, where $\beta_{y}^{T}$ and $\beta_{y}^{C}$ are the vertical temperature and composition gradients, respectively. The dimensionless equations for $\bm{u}$, $\bm{b}$, $\Theta$ and $\gamma$ are given by,
$E_{\eta}\bigg(\frac{\partial\bm{u}}{\partial t}+u_{c}\,z\frac{\partial\bm{u}}{\partial x}+u_{c}u_{z}\hat{\bm{e}}_{x}\bigg)+\hat{\bm{e}}_{z}\times\bm{u}=-\nabla p^{\star}+(\nabla\times\bm{b})\times\bm{B}_{0}$
$+Ra_{\ell,V}^{T}\Theta\hat{\bm{e}}_{y}+Ra_{\ell,V}^{C}\gamma\hat{\bm{e}}_{y}+E\nabla^{2}\bm{u},$ (58)
$\frac{\partial\bm{b}}{\partial t}+u_{c}\,z\frac{\partial\bm{b}}{\partial x}=u_{c}b_{z}\hat{\bm{e}}_{x}+(\bm{B}_{0}\cdot\nabla)\bm{u}+\eta\nabla^{2}\bm{b},$ (59)
$\frac{\partial\Theta}{\partial t}+u_{c}z\frac{\partial\Theta}{\partial x}+u_{y}+u_{z}\beta^{\star}=q^{T}\nabla^{2}\Theta,$ (60)
$\frac{\partial\gamma}{\partial t}+u_{c}z\frac{\partial\gamma}{\partial x}+u_{y}=q^{C}\nabla^{2}\gamma,$ (61)
$\nabla\cdot\bm{u}=\nabla\cdot\bm{b}=0.$ (62)
The dimensionless parameters in (58), based on the length scale of the density disturbance, are the Ekman number $E=\nu/2\varOmega\delta^{2}$, magnetic Ekman number $E_{\eta}=\eta/2\varOmega\delta^{2}$, and local vertical compositional and thermal Rayleigh numbers, $Ra_{\ell,V}^{C}=g\alpha^{C}|\beta_{y}^{C}|\delta^{2}/2\varOmega\eta$ and $Ra_{\ell,V}^{T}=g\alpha^{T}|\beta_{y}^{T}|\delta^{2}/2\varOmega\eta$, respectively. Here, $\alpha^{T}$ and $\alpha^{C}$ are the thermal and compositional expansion coefficients, respectively. Additionally, $q^{T}=\kappa^{T}/\eta$, $q^{C}=\kappa^{C}/\eta$ and $\beta^{\star}=\beta_{z}/\beta_{y}^{C}$, where $\kappa_{T}$ and $\kappa_{C}$ are the diffusivities of temperature and composition and $\beta_{z}$ is the horizontal temperature gradient.
For the rapidly rotating system considered in §2, the limit $\nu,\kappa_{C},\kappa_{T}\ll\eta$ results in $E,q^{C},q^{T}\to 0$ while $E_{\eta}\ll 1$ retains a small but finite value. In this limit, a plane wave solution of the form $\hat{u}_{z}\sim\mbox{e}^{\mathrm{i}\lambda t}$ gives the characteristic equation,
$\lambda^{5}$ $-2\mathrm{i}\omega_{\eta}\lambda^{4}-(\omega_{C}^{2}+2\omega_{M}^{2}+{\omega^{C}_{A,V}}^{2}+{\omega^{T}_{A,V}}^{2}-2{\omega_{A,H}^{2}}+\omega_{\eta}^{2})\lambda^{3}$ (63)
$+2\mathrm{i}\omega_{\eta}(\omega_{C}^{2}+\omega_{M}^{2}+{\omega^{C}_{A,V}}^{2}+{\omega^{T}_{A,V}}^{2}-2{\omega_{A,H}^{2}})\lambda^{2}$
$+\big(\omega_{C}^{2}\omega_{\eta}^{2}+\omega_{M}^{4}+({\omega^{C}_{A,V}}^{2}+{\omega^{T}_{A,V}}^{2}-2{\omega_{A,H}^{2}})(\omega_{M}^{2}+\omega_{\eta}^{2})\big)\lambda$
$-\mathrm{i}\omega_{\eta}\omega_{M}^{2}({\omega^{C}_{A,V}}^{2}+{\omega^{T}_{A,V}}^{2}-{\omega_{A,H}^{2}})=0,$
where the dimensionless fundamental frequencies are,
$\begin{aligned} &\omega_{C}^{2}=\frac{1}{E_{\eta}^{2}}\frac{k_{z}^{2}}{k^{2}},\quad\omega_{M}^{2}=\frac{(\bm{B}\cdot\bm{k})^{2}}{E_{\eta}},\quad{\omega_{A,V}^{T}}^{2}=\frac{Ra_{\ell,V}^{T}}{E_{\eta}}\frac{k_{z}^{2}}{k^{2}},\quad{\omega_{A,V}^{C}}^{2}=\frac{Ra_{\ell,V}^{C}}{E_{\eta}}\frac{k_{z}^{2}}{k^{2}},\quad\\ &\omega_{A,H}^{2}=\frac{Ra_{\ell,H}}{2E_{\eta}}\frac{k_{z}^{2}}{k^{2}},\quad\omega_{\eta}^{2}=k^{4}.\end{aligned}$ (64a–f)
The local horizontal Rayleigh number $Ra_{\ell,H}$ has the same definition as in (34). For the inequality $|\omega_{C}|\gg|\omega_{M}|\gg|\omega_{A,V}^{C}|,\,|\omega_{A,V}^{T}|,\,|\omega_{A,H}|\gg|\omega_{\eta}|$, the roots of the characteristic equation (63) are approximated by,
$\lambda_{1,2}$ $\approx$ $\pm\bigg(\omega_{C}+\frac{\omega_{M}^{2}}{\omega_{C}}\bigg)+\mathrm{i}\,\frac{\omega_{M}^{2}\omega_{\eta}}{\omega_{C}^{2}},$ (65)
$\lambda_{3,4}$ $\approx$ $\pm\bigg(\frac{\omega_{M}^{2}}{\omega_{C}}+\frac{{\omega_{A,V}^{C}}^{2}+{\omega_{A,V}^{T}}^{2}-2\omega_{A,H}^{2}}{2\omega_{C}}\bigg)+\mathrm{i}\,\omega_{\eta}\,\bigg(1-\frac{{\omega_{A,V}^{C}}^{2}+{\omega_{A,V}^{T}}^{2}-\omega_{A,H}^{2}}{2\omega_{M}^{2}}\bigg),$ (66)
$\lambda_{5}$ $\approx$ $\mathrm{i}\,\omega_{\eta}\frac{{\omega_{A,V}^{C}}^{2}+{\omega_{A,V}^{T}}^{2}-\omega_{A,H}^{2}}{\omega_{M}^{2}},$ (67)
Here, $\lambda_{1,2}$ and $\lambda_{3,4}$ give the approximate frequencies of the oppositely travelling fast ($\omega_{f}$) and slow ($\omega_{s}$) MAC waves, respectively. From (66), the resultant buoyancy frequency is given by
$\omega_{A}^{2}={\omega_{A,V}^{C}}^{2}+{\omega_{A,V}^{T}}^{2}-2\omega_{A,H}^{2},$ (68)
which indicates the complementarity between the vertical buoyancies of composition and temperature and the horizontal buoyancy of temperature. The resultant local Rayleigh number is then given by
$Ra_{\ell}=\frac{g\big|\alpha^{C}\beta_{y}^{C}+\alpha^{T}\beta_{y}^{T}-2\alpha^{T}\beta_{z}k_{y}/k_{z}\big|\delta^{2}}{2\varOmega\eta},$ (69)
The solutions for $\hat{u}_{z}$, $\hat{\Theta}$ and $\hat{\gamma}$ are,
$\big[\hat{u}_{z},\hat{\Theta},\hat{\gamma}\big]=\sum_{m=1}^{5}\big[D_{m},G_{m},Q_{m}\big]\mbox{e}^{\mathrm{i}\lambda_{m}t},$ (70)
where the coefficients $D_{m}$, $G_{m}$ and $Q_{m}$ are evaluated from the initial conditions for $\hat{u}_{z}$, $\hat{\Theta}$, $\hat{\gamma}$ and their time derivatives. For instance, the initial conditions for $\hat{u}_{z}$ and its time derivatives are given by
$\mathrm{i}^{n}\sum_{m=1}^{5}D_{m}\lambda_{m}^{n}=\bigg(\frac{\partial^{n}\hat{u}_{z}}{\partial t^{n}}\bigg)_{t=0}=d_{n+1},\ n=0,1,2,3,4,$ (71)
where
$d_{1}$ $=\hat{u}_{z}\big|_{t=0}=0,$ (72)
$d_{2}$ $=\frac{\partial{\hat{u}_{z}}}{\partial{t}}\big|_{t=0}=\bigg({\omega_{A,V}^{C}}^{2}\hat{\Theta}_{0}+{\omega_{A,V}^{T}}^{2}\hat{\gamma}_{0}\bigg)\frac{k_{y}}{k_{z}},$
$d_{3}$ $=\frac{\partial^{2}\hat{u}_{z}}{\partial{t^{2}}}\big|_{t=0}=0,$
$d_{4}$ $=\frac{\partial^{3}{\hat{u}_{z}}}{\partial{t^{3}}}\big|_{t=0}=-d_{2}(\omega_{M}^{2}+\omega_{C}^{2}+\omega_{A}^{2}),$
$d_{5}$ $=\frac{\partial^{4}{\hat{u}_{z}}}{\partial{t^{4}}}\big|_{t=0}=d_{2}\omega_{\eta}\omega_{M}^{2},$
where $\hat{\Theta}_{0}$ and $\hat{\gamma}_{0}$ are the initial perturbations of temperature and composition respectively, both of which have the Gaussian distribution as in (2) in Cartesian coordinates. The coefficients of the fast and slow wave parts of $\hat{u}_{z}$ are then obtained from the roots of (63), as follows:
$D_{1}$ $=\dfrac{d_{5}-\mathrm{i}\,d_{4}(\lambda_{2}+\lambda_{3}+\lambda_{4}+\lambda_{5})+\mathrm{i}\,d_{2}(\lambda_{2}\lambda_{4}\lambda_{5}+\lambda_{3}\lambda_{4}\lambda_{5}+\lambda_{2}\lambda_{3}\lambda_{4}+\lambda_{2}\lambda_{3}\lambda_{5})}{(\lambda_{1}-\lambda_{2})(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{4})(\lambda_{1}-\lambda_{5})},$ (73)
$D_{3}$ $=\dfrac{d_{5}-\mathrm{i}\,d_{4}(\lambda_{1}+\lambda_{2}+\lambda_{4}+\lambda_{5})+\mathrm{i}\,d_{2}(\lambda_{1}\lambda_{4}\lambda_{5}+\lambda_{2}\lambda_{4}\lambda_{5}+\lambda_{1}\lambda_{2}\lambda_{4}+\lambda_{1}\lambda_{2}\lambda_{5})}{(\lambda_{3}-\lambda_{1})(\lambda_{3}-\lambda_{2})(\lambda_{3}-\lambda_{4})(\lambda_{3}-\lambda_{5})},$ (74)
Further, $\hat{u}_{y}$ is calculated from equation (18). The spectral coefficients of $\hat{\Theta}$ and $\hat{\gamma}$ are obtained in a similar way (Appendix B).
In two-component convection, both thermal and compositional buoyancies drive the convection and are responsible for magnetic field generation. The fraction of thermal power relative to the total convective power is given by,
$f^{T}=\frac{P^{T}}{P^{T}+P^{C}}\times 100\%,$ (75)
where $P^{T}$ and $P^{C}$ are the dimensionless thermal and compositional powers respectively, defined by
$P^{T}=Ra_{\ell,V}^{T}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\bm{u}\cdot\Theta\hat{\bm{e}}_{y}\,\mbox{d}y\mbox{d}z,\quad P^{C}=Ra_{\ell,V}^{C}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\bm{u}\cdot\gamma\hat{\bm{e}}_{y}\,\mbox{d}y\mbox{d}z.$ (76)
The integrals in (76) are computed at $x=0$ for the limits $\pm 20$ in $(y,z)$ (The integrals of $u_{y}\Delta T_{0}$ and $u_{y}\Delta C_{0}$ vanish).
4.2 Complementarity of vertical and horizontal buoyancies in two-component convection
Figure 16: Variation of the squares of frequencies with relative horizontal buoyancy in two-component magnetoconvection. The dotted vertical line in (a) corresponds to $|\omega_{A}|\approx|\omega_{M}|$, when the slow wave frequency $\omega_{s}$ goes to zero. Panel (b) gives the decomposition of $\omega_{A}^{2}$ into its three parts consisting of the vertical buoyancies of composition and temperature (${\omega^{C}_{A,V}}^{2}$, ${\omega^{T}_{A,V}}^{2}$) and the horizontal buoyancy of temperature ($\omega_{A,H}^{2}$). The parameters used are $E_{\eta}=2\times 10^{-5}$, $B^{2}_{peak}=200$ and $f^{T}=10$% at time $t/t_{\eta}=10^{-2}$.
In §2.4, the complementarity between vertical and horizontal buoyancies at the state of vanishing slow MAC waves was noted. In two-component convection, a complementarity exists between the vertical buoyancies of composition and temperature and the horizontal buoyancy of temperature in suppressing the slow MAC waves (figure 16). Figure 16(a) shows the variation of the squares of the fundamental frequencies and the slow MAC wave frequency with relative horizontal buoyancy, whereas figure 16(b) presents the decomposition of $\omega_{A}^{2}$ into its three constituent parts. While the compositional Rayleigh number $Ra^{C}_{\ell,V}$ is held constant (see below), the thermal Rayleigh number is determined by the power ratio $f^{T}$, given by (75). In figure 16(a), $f^{T}$ is set to 10% and the horizontal buoyancy is progressively increased. When the resultant $\omega_{A}$ obtained from (68) matches $\omega_{M}$, the slow MAC waves disappear.
The analysis in figure 16 is then performed for a range of $f^{T}$, and the results are summarized in table 4.
4.2.1 A constraint on the lateral heat flux variation in Earth’s lower mantle
In one-component convective dynamos, the heterogeneity factor $q^{*}$ that causes the suppression of slow MAC waves may take values of O(1)O(1)$O(10)$; see figure 15(d). In two-component convection, we aim to place an upper bound on $q^{*}$ for Earth using plausible values of the peak magnetic field intensity and the Rayleigh number in the core. For the dipole-dominated regime given by $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|$, the square of the peak value of the scaled magnetic field should take values of $O(10^{2})$ (Sreenivasan and Maurya, 2021; Varma and Sreenivasan, 2022) while the mean square value of the field could be $O(1)$, as suggested by observations (Gillet et al., 2010). Furthermore, if strong compositional buoyancy in the core by itself maintains the dynamo in a dipolar state below the threshold for polarity transitions, a value of $Ra^{C}_{\ell,V}$ of $O(10^{3})$ is reasonable (Majumder et al., 2024). For Earth, this value of the local Rayleigh number corresponds to a vertical Rayleigh number $O(10^{3})$ times the critical Rayleigh number for onset of nonmagnetic convection, essential for reproducing the observed polar circulation of 0.6–0.9 yr-1 (Hulot et al., 2002) in the low-inertia geodynamo (see Majumder and Sreenivasan, 2023). Now, using the complementarity of the buoyancy frequencies (68), the value of the relative horizontal buoyancy $Ra_{\ell,H}/Ra^{T}_{\ell}$ for polarity transitions based on the condition of vanishing slow MAC waves, $|\omega_{A}|\approx|\omega_{M}|$, may be obtained for a range of thermal power fractions $f^{T}$. Since thermal buoyancy is not expected to contribute more than 25% of the total buoyancy power for Earth’s core (Lister and Buffett, 1995), the geodynamo can exist below the threshold for polarity reversals for $Ra_{\ell,H}/Ra^{T}_{\ell}>0.5$ (table 4), which must in turn correspond to large lower-mantle heat flux variations $q^{*}$ of $O(10)$ as per figure 15(d). The self-consistent two-component dynamo model, subject to heterogeneous outer boundary heat flux derived from the seismic shear wave velocity in Earth’s lower mantle, further substantiates this point (see table 5 in Appendix C). The estimated value of $q^{*}$ is regarded as an upper bound because it corresponds to the suppression of the slow MAC waves, at which state the dynamo exhibits reversals consistent with the occasional reversals observed in Earth. Beyond this limit, the dynamo transitions to a multipolar state. For relatively small horizontal buoyancy, the dynamo would be comfortably placed in the dipole-dominated regime without polarity transitions. In short, for nearly invariant vertical Rayleigh numbers of composition and temperature, the magnitude of the horizontal Rayleigh number would determine whether the dynamo operates in a non-reversing dipolar state, a reversing state or a multipolar state.
SL $B^{2}_{peak}$ $f^{T}\%$ $Ra^{C}_{\ell,V}$ $Ra^{T}_{\ell,V}$ $Ra_{\ell,H}$ $\omega_{A,V}^{2}/\omega_{M}^{2}$ $\omega_{A,H}^{2}/\omega_{M}^{2}$ $Ra_{\ell,H}/Ra^{T}_{\ell}$
1 200 10 2500 875 2461 0.57 0.215 0.74
2 200 12 2500 966 2370 0.59 0.207 0.71
3 200 15 2500 1045 2291 0.61 0.195 0.64
4 200 20 2500 1230 2106 0.63 0.185 0.62
5 200 25 2500 1637 1699 0.68 0.160 0.51
6 200 30 2500 1791 1545 0.72 0.140 0.46
7 200 40 2500 2098 1238 0.78 0.110 0.37
8 200 50 2500 2500 836 0.83 0.085 0.25
Table 4: Calculation of the relative intensity of horizontal buoyancy (last column) in two-component linear magnetoconvection for states where the slow MAC waves disappear. The thermal power ratio $f^{T}$ is defined in (75); $Ra^{C}_{\ell,V}$, $Ra^{T}_{\ell,V}$ are the local vertical compositional and thermal Rayleigh numbers respectively. In addition, $Ra_{\ell,H}$ is the local horizontal Rayleigh number. The resultant local thermal Rayleigh number is $Ra_{\ell}^{T}=Ra^{T}_{\ell,V}+Ra_{\ell,H}$. The parameters are $E_{\eta}=2\times 10^{-5}$ and $t=0.01$.
5 Concluding remarks
The present study investigates the dipole–multipole transition through the analysis of MHD wave motions in rapidly rotating dynamos subject to inhomogeneous heat flux at the outer boundary. The regime in focus is that of low inertia, where the nonlinear inertial force is small relative to the Coriolis force not only on the length scale of the planetary core but also on the characteristic length scale of convection. The dipole–multipole transition corresponds to the approximate parity between $|\omega_{M}|$ and $|\omega_{A}|$, where $|\omega_{M}|$ is based on the peak field intensity and $|\omega_{A}|$ is a resultant buoyancy frequency derived from the vertical and horizontal (lateral) buoyancy frequencies. The present study focuses on non-axisymmetric outer boundary heat flux, motivated by mantle convection models. That said, as indicated by earlier studies (e.g. Glatzmaier et al., 1999), a high equatorial axisymmetric boundary heat flux can also induce the dipole–multipole transition in rotating dynamos. An axisymmetric positive equatorial heat flux anomaly is analogous to the enhanced equatorial heat flux in a model with homogeneous boundary heat flux (Majumder et al., 2024), where the parity between $|\omega_{M}|$ and $|\omega_{A}|$ in equation (1) causes the disappearance of the slow waves, and in turn, the polarity transition. For the axisymmetric, equatorially anti-symmetric $Y_{1}^{0}$ heterogeneity, the resultant $\beta$ in equation (49) must be evaluated with $\beta_{z}$ taken under a modulus sign in regions where $\partial T/\partial r<0$, such that an increase in $q^{*}$ (or, equivalently, in the magnitude of $\beta_{z}$) leads to an increase in the resultant buoyancy frequency $|\omega_{A}|$. Under this procedure, preferential cooling of either the Northern or the Southern hemisphere has the same outcome, likely inducing polarity transitions. However, the theoretical model in §2 does not predict the total suppression of convection under stable stratification, the analysis of which should be the subject of a future study.
A dipole-dominated, strong-field dynamo with a mean square magnetic field intensity $B^{2}/(2\varOmega\rho\mu\eta)=O(1)$ undergoes a polarity transition through the progressive increase of horizontal buoyancy induced by an equatorially anti-symmetric heat flux heterogeneity at the outer boundary. The transition occurs not by the loss of equatorial symmetry of the convection columns, but by the selective suppression of the slow MAC waves. Polarity reversals lie in a range of horizontal buoyancies between the dipolar and multipolar regimes. A non-axisymmetric, equatorially symmetric heat flux variation does not induce the transition even at high values of the dimensionless heterogeneity $q^{*}$ since the mean temperature gradient at the equator is unaffected for a fixed vertical buoyancy. A boundary heterogeneity consisting of comparable magnitudes of symmetric and anti-symmetric variations induces the polarity transition at a value of $q^{*}$ which is of the same order as that for the transition induced by a purely anti-symmetric variation. Whether the rapid collapse of the axial dipole during reversals may be correlated with the decay of the slow MAC waves when $|\omega_{A}|\approx|\omega_{M}|$ is a problem that requires further investigation.
The present study proposes a dynamical constraint on the upper bound of the heat flux heterogeneity based on the complementarity between the vertical and horizontal buoyancies in suppressing the slow waves, noted in both the linear magnetoconvection and nonlinear dynamo models. In a single-component (thermal) buoyancy-driven dynamo, a dipole solution may be possible for any value in the range 0–0.75 of the relative horizontal buoyancy, $Ra_{\ell,H}/Ra_{\ell}$ (figure 15(c)). However, the fact that compositional buoyancy is dominant in the two-component geodynamo, together with the known order of magnitude of the peak field intensity in the inertia-free limit, suggest $Ra_{\ell,H}/Ra_{\ell}>0.5$, corresponding to $q^{*}$ of $O(10)$ at which an axial dipole can exist. From the point of view of global mantle convection (Olson et al., 2015) and the regional stratification of the outer core (Mound et al., 2019), the maximum variation of heat flux at Earth’s CMB could be higher than the mean superadiabatic heat flux.
For comparable thermal power ratios $f^{T}$, the magnitudes of the upper bound of $q^{*}$ predicted by the two-component magnetoconvection model (§4.1) and obtained from the self-consistent two-component dynamo simulations (Appendix C) are comparable. A lower bound for the heterogeneity may also be obtained from dynamo simulations, based on the minimum variation needed to obtain the hemispherical (east–west) variability in the high-latitude magnetic flux in Earth (Sahoo and Sreenivasan, 2020) and the mantle-induced heat flux heterogeneity at the inner core boundary (Sreenivasan and Gubbins, 2011).
In §4.1, the chosen value of the local compositional Rayleigh number $Ra_{\ell,V}^{C}$ is well supported by observations of the polar core flow (Majumder and Sreenivasan, 2023). That said, the value of $Ra_{\ell,V}^{C}$ in Earth’s core may have varied in the past. Based on the complementarity of buoyancy frequencies in (68), small vertical Rayleigh numbers require large horizontal Rayleigh numbers, and in turn, large $q^{*}$ to achieve reversals. For the present-day Earth, with nearly invariant vertical Rayleigh numbers of composition and temperature, values of the relative horizontal buoyancy much lower than that needed to suppress the slow MAC waves in the core would place the dynamo in a non-reversing dipolar state. This provides a way to explain the existence of geomagnetic superchrons, the long periods in Earth’s history without polarity reversals. That said, even for relative horizontal buoyancies >0.5>0.5, the ratio of symmetric to anti-symmetric heterogeneity in heat flux at Earth’s CMB can vary in geological time. For example, Courtillot and Olson (2007) suggest a relation between superchrons and the formation of mantle plumes, which control the heat flow across the CMB. If a plume forms near the equator, the heat flux at the CMB is dominated by an equatorially symmetric heat flux pattern, promoting a dipole-dominated non-reversing state. On the other hand, if a plume forms far from the equator, the resulting anti-symmetric heat flux pattern suppresses the slow MAC waves and produces a regime conducive to reversals.
Acknowledgments
This study was supported in part by Research Grant MoE-STARS/STARS-1/504 under Scheme for Transformational and Advanced Research in Sciences awarded by the Ministry of Education (India) and in part by Research grant CRG/2021/002486 awarded by the Science and Engineering Research Board (India). The computations were performed on Param Pravega, the supercomputer at the Indian Institute of Science, Bangalore.
Declaration of interests
The authors report no conflict of interest.
Appendix A Dynamo with outer boundary heat flux derived from seismic tomography
A dynamo subject to an outer boundary heat flux varying linearly as the seismic shear wave velocity in the lower mantle (Masters et al., 1996) produces polarity reversals at sufficiently large heat flux heterogeneity, as shown by the evolution of the axial dipole colatitude in figure 17. This heat flux heterogeneity is a combination of equatorially symmetric and anti-symmetric variations of comparable magnitude. The resultant basic state (mean) temperature gradient $\beta$ is given by (49), wherein $\beta_{s}$ measures the basic state gradient with homogeneous boundary heat flux and the non-zero $\beta_{z}$ at the equator measures the anti-symmetric part of the variation. The square of the resultant buoyancy frequency $\omega_{A}$, shown in figures 18 (a) & (c) at a section $z=0.2$ below the equatorial plane, is proportional to the resultant gradient $\beta$. In the non-reversing dipolar run at $q^{*}=10$, $|\omega_{M}|^{2}>|\omega_{A}|^{2}$ in several regions (figure 18c), suggesting the generation of slow MAC waves. However, in the reversing run at $q^{*}=13$, $|\omega_{M}|^{2}\approx|\omega_{A}|^{2}$ in this section (figure 18d), suggesting the suppression of the slow waves.
Figure 17: Evolution of dipole colatitude with time (measured in units of the magnetic diffusion time) for the dynamo subject to heterogeneous outer boundary heat flux based on the seismic shear wave velocity in Earth’s lower mantle at $q^{*}=13$. The other parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
(a)               (b)
Refer to caption Refer to caption
(c)               (d)
Refer to caption Refer to caption
Figure 18: Section plots at height $z=0.2$ below the equator showing $\omega_{A}^{2}$ (a, c) and $|\omega_{M}^{2}/\omega_{A}^{2}|$ (b, d) for the composite outer boundary heat flux heterogeneity based on the seismic shear wave velocity in Earth’s lowermost mantle. Two values of the heterogeneity are considered, $q^{*}=10$ (a, b) and $q^{*}=13$ (c, d). The other parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
Appendix B Spectral coefficients for the perturbations of temperature and composition
From (70), the initial conditions for $\hat{\Theta}$ and its time derivatives are given by
$\mathrm{i}^{n}\sum_{m=1}^{5}G_{m}\lambda_{m}^{n}=\bigg(\frac{\partial^{n}\hat{\Theta}}{\partial t^{n}}\bigg)_{t=0}=g_{n+1},\ n=0,1,2,3,4,$ (77)
where
$g_{1}$ $=\hat{\Theta}\big|_{t=0}=\hat{\Theta}_{0},$ (78)
$g_{2}$ $=\frac{\partial{\hat{\Theta}}}{\partial{t}}\big|_{t=0}=0,$
$g_{3}$ $=\frac{\partial^{2}\hat{\Theta}}{\partial{t^{2}}}\big|_{t=0}=a_{2}\bigg(\beta_{y}^{T}\frac{k_{y}}{k_{z}}-\beta_{z}\bigg),$
$g_{4}$ $=\frac{\partial^{3}\hat{\Theta}}{\partial{t^{3}}}\big|_{t=0}=a_{3}\bigg(\beta_{y}^{T}\frac{k_{y}}{k_{z}}-\beta_{z}\bigg),$
$g_{5}$ $=\frac{\partial^{4}\hat{\Theta}}{\partial{t^{4}}}\big|_{t=0}=a_{4}\bigg(\beta_{y}^{T}\frac{k_{y}}{k_{z}}-\beta_{z}\bigg).$
Here, the coefficients $a_{2}$, $a_{3}$, $a_{4}$ are defined in (24).
The coefficients of the fast and slow wave components of $\hat{\Theta}$ are obtained using the roots of equation (63). For example,
$G_{1}$ $=\dfrac{g_{5}-\mathrm{i}\,g_{4}(\lambda_{2}+\lambda_{3}+\lambda_{4}+\lambda_{5})+\mathrm{i}\,g_{2}(\lambda_{2}\lambda_{4}\lambda_{5}+\lambda_{3}\lambda_{4}\lambda_{5}+\lambda_{2}\lambda_{3}\lambda_{4}+\lambda_{2}\lambda_{3}\lambda_{5})}{(\lambda_{1}-\lambda_{2})(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{4})(\lambda_{1}-\lambda_{5})},$ (79)
$G_{3}$ $=\dfrac{g_{5}-\mathrm{i}\,g_{4}(\lambda_{1}+\lambda_{2}+\lambda_{4}+\lambda_{5})+\mathrm{i}\,g_{2}(\lambda_{1}\lambda_{4}\lambda_{5}+\lambda_{2}\lambda_{4}\lambda_{5}+\lambda_{1}\lambda_{2}\lambda_{4}+\lambda_{1}\lambda_{2}\lambda_{5})}{(\lambda_{3}-\lambda_{1})(\lambda_{3}-\lambda_{2})(\lambda_{3}-\lambda_{4})(\lambda_{3}-\lambda_{5})},$ (80)
From (70), the initial conditions for $\hat{\gamma}$ and its time derivatives are given by
$\mathrm{i}^{n}\sum_{m=1}^{5}Q_{m}\lambda_{m}^{n}=\bigg(\frac{\partial^{n}\hat{\gamma}}{\partial t^{n}}\bigg)_{t=0}=q_{n+1},\ n=0,1,2,3,4,$ (81)
where
$q_{1}$ $=\hat{\gamma}\big|_{t=0}=\hat{\gamma}_{0},$ (82)
$q_{2}$ $=\frac{\partial{\hat{\gamma}}}{\partial{t}}\big|_{t=0}=0,$
$q_{3}$ $=\frac{\partial^{2}\hat{\gamma}}{\partial{t^{2}}}\big|_{t=0}=a_{2}\beta_{y}^{C}\frac{k_{y}}{k_{z}},$
$q_{4}$ $=\frac{\partial^{3}\hat{\gamma}}{\partial{t^{3}}}\big|_{t=0}=a_{3}\beta_{y}^{C}\frac{k_{y}}{k_{z}},$
$q_{5}$ $=\frac{\partial^{4}\hat{\gamma}}{\partial{t^{4}}}\big|_{t=0}=a_{4}\beta_{y}^{C}\frac{k_{y}}{k_{z}}.$
The coefficients of the fast and slow wave components of $\hat{\gamma}$ are obtained using the roots of equation (63). For example,
$Q_{1}$ $=\dfrac{q_{5}-\mathrm{i}\,q_{4}(\lambda_{2}+\lambda_{3}+\lambda_{4}+\lambda_{5})+\mathrm{i}\,q_{2}(\lambda_{2}\lambda_{4}\lambda_{5}+\lambda_{3}\lambda_{4}\lambda_{5}+\lambda_{2}\lambda_{3}\lambda_{4}+\lambda_{2}\lambda_{3}\lambda_{5})}{(\lambda_{1}-\lambda_{2})(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{4})(\lambda_{1}-\lambda_{5})},$ (83)
$Q_{3}$ $=\dfrac{q_{5}-\mathrm{i}\,q_{4}(\lambda_{1}+\lambda_{2}+\lambda_{4}+\lambda_{5})+\mathrm{i}\,q_{2}(\lambda_{1}\lambda_{4}\lambda_{5}+\lambda_{2}\lambda_{4}\lambda_{5}+\lambda_{1}\lambda_{2}\lambda_{4}+\lambda_{1}\lambda_{2}\lambda_{5})}{(\lambda_{3}-\lambda_{1})(\lambda_{3}-\lambda_{2})(\lambda_{3}-\lambda_{4})(\lambda_{3}-\lambda_{5})},$ (84)
Appendix C Two-component nonlinear dynamo model
(a)                  (b)
Refer to caption Refer to caption
Figure 19: Evolution of dipole colatitude with magnetic diffusion time for (a) $q^{*}=10,19$ (stable dipolar) and (b) $q^{*}=20$ (reversing) for the dynamo subject to heterogeneous outer boundary heat flux based on the seismic shear wave velocity in Earth’s lower mantle. (c) Snapshot of the radial magnetic field at the outer boundary for $q^{*}=10$. The dynamo parameters are $E=6\times 10^{-5},Pm=5,Sc=5,Pr=0.5,Ra^{T}=130,Ra^{C}=18000$.
$q^{*}$ $N_{r}$ $l_{max}$ RmRm $Ro_{l}$ $l_{E}$ $l_{C}$ $\bar{m}$ $\bar{k}_{s}$ $\bar{k}_{z}$ $E_{k}$ $E_{m}$ $B^{2}_{peak}$ $B^{2}_{rms}$ $f_{dip}$ $\dfrac{Ra_{\ell,H}}{Ra_{\ell}^{T}}$ Type
$\times 10^{5}$ $\times 10^{5}$ D/R/M
0 144 144 450 0.025 24 15 7 4.23 4.47 14.68 25.03 223 4.36 0.63 0.00 D
1 144 144 455 0.026 25 15 7 4.39 4.45 17.97 22.52 216 4.15 0.62 0.07 D
3 144 144 461 0.027 25 16 7 4.46 4.44 15.08 18.32 211 2.79 0.62 0.19 D
5 144 144 464 0.027 25 16 7 4.32 4.37 15.12 18.41 205 2.85 0.61 0.28 D
7 144 144 469 0.027 26 16 7 4.38 4.43 15.25 18.53 202 2.83 0.61 0.35 D
10 144 144 470 0.028 28 16 7 4.48 4.33 15.55 18.84 190 2.98 0.59 0.43 D
12 144 144 472 0.031 29 17 7 4.45 4.43 16.31 16.95 186 2.54 0.57 0.47 D
15 144 144 473 0.031 30 17 7 4.49 4.30 16.43 17.86 183 2.57 0.53 0.51 D
18 144 144 482 0.031 31 17 7 4.43 4.35 16.49 17.02 179 2.63 0.47 0.57 D
19 144 144 496 0.032 31 17 7 4.30 4.31 16.29 16.87 175 3.42 0.46 0.58 D
20 144 144 505 0.032 32 17 7 4.46 4.33 18.46 8.87 136 1.09 0.31 0.60 R
25 144 144 516 0.032 33 17 7 4.48 4.46 18.74 8.42 107 1.08 0.30 0.62 R
Table 5: Summary of the main input and output parameters of the two-component dynamo simulations considered in this study. Here, $q^{*}$ is the dimensionless measure of the boundary heterogeneity, defined in (45), $N_{r}$ is the number of radial grid points, $l_{max}$ is the maximum spherical harmonic degree, RmRm is the magnetic Reynolds number, $Ro_{\ell}$ is the local Rossby number, $l_{C}$ and $l_{E}$ are the mean spherical harmonic degrees of convection and energy injection, $\bar{m}$ is the mean spherical harmonic order in the range $l\leq l_{E}$ , $\bar{k}_{s}$ and $\bar{k}_{z}$ are the mean $s$ and $z$ wavenumbers in the range $l\leq l_{E}$, $E_{k}$ and $E_{m}$ are the time-averaged total kinetic and magnetic energies, $f_{dip}$ is the relative axial dipole field strength, $B^{2}_{peak}$ is the square of the peak field in the saturated dynamo, and $B^{2}_{rms}$ is the measured mean square value of the field in the spherical shell. Type ‘D’ and ‘R’ denotes dipolar and reversing dynamos, respectively. The local horizontal Rayleigh number is given by $Ra_{\ell,H}$ and the local thermal Rayleigh number is $Ra_{\ell}^{T}$. The dynamo parameters are $E=6\times 10^{-5},Pm=5,Sc=5,Pr=0.5,Ra^{T}=130,Ra^{C}=18000$ and the thermal power ratio $f^{T}=11\%$, where $Ra^{T}$ and $Ra^{C}$ are the modified thermal and compositional Rayleigh numbers. Here, $Ra^{T}$ is set to its critical value for the onset of nonmagnetic convection with homogeneous boundary heat flux while $Ra^{C}$ is $\approx 900\times$ its critical value for the onset of convection.
A thermochemically driven dynamo is considered within an electrically conducting fluid, which is confined between two concentric, co-rotating spherical surfaces corresponding to the inner core boundary (ICB) and the core–mantle boundary (CMB). The ratio of the inner radius $r_{i}$ to the outer radius $r_{o}$ is set to 0.35. Lengths are scaled by $L$ and time is scaled by $L^{2}/\eta$, where $\eta$ is the magnetic diffusivity. The velocity is scaled by $\eta/L$ and the magnetic field is scaled by $(2\varOmega\mu\eta\rho)^{1/2}$, where $\varOmega$ is the angular velocity of rotation, $\mu$ is the magnetic permeability and $\rho$ is the fluid density. The temperature is scaled by $\beta^{T}L$ and the composition is scaled by $\beta^{C}L$, where $\beta^{T}$ and $\beta^{C}$ are the mean thermal and compositional gradients within the shell, respectively. In the Boussinesq approximation, the governing non-dimensional magnetohydrodynamic (MHD) equations for velocity, magnetic field, temperature, and composition are as follows:
$EPm^{-1}\Bigl(\frac{\partial{\bm{u}}}{\partial t}+(\nabla\times{\bm{u}})\times{\bm{u}}\Bigr)+{\hat{\bm{z}}}\times{\bm{u}}=-\nabla p^{\star}+Ra^{T}\,T\,{\bm{r}}\,$
$+Ra^{C}\,C\,{\bm{r}}+(\nabla\times{\bm{B}})\times{\bm{B}}+E\nabla^{2}{\bm{u}},$ (85)
$\frac{\partial{\bm{B}}}{\partial t}=\nabla\times({\bm{u}}\times{\bm{B}})+\nabla^{2}{\bm{B}},$ (86)
$\frac{\partial T}{\partial t}+({\bm{u}}\cdot\nabla)T=PmPr^{-1}\,\nabla^{2}T+S_{o},$ (87)
$\frac{\partial C}{\partial t}+({\bm{u}}\cdot\nabla)C=PmSc^{-1}\,\nabla^{2}C+S_{i},$ (88)
$\nabla\cdot{\bm{u}}=\nabla\cdot{\bm{B}}=0.$ (89)
The modified pressure, $p^{*}$ is expressed as $p+\frac{1}{2}E\,Pm^{-1}\,|\bm{u}|^{2}$. The dimensionless parameters governing the system are the Ekman number $E=\nu/2\varOmega L^{2}$, the Prandtl number $Pr=\nu/\kappa^{T}$, the Schmidt number $Sc=\nu/\kappa^{C}$ and the magnetic Prandtl number, $Pm=\nu/\eta$. The modified thermal and compositional Rayleigh numbers are given by $Ra^{T}=g\alpha^{T}\beta^{T}L^{2}/2\Omega\eta$ and $Ra^{C}=g\alpha^{C}\beta^{C}L^{2}/2\Omega\eta$, respectively. Here, $g$ is the gravitational acceleration, $\nu$ is the kinematic viscosity, $\kappa^{T}$ and $\kappa^{C}$ are the thermal and compositional diffusivities, and $\alpha^{T}$ and $\alpha^{C}$ are the coefficients of thermal and compositional expansion, respectively.
No-slip and electrically insulating conditions are imposed at the two boundaries. The basic state temperature profile consists of a combination of basal and internal heating,
$\frac{\partial T_{0}}{\partial r}=-\frac{A}{r^{2}}-B\,r,$ (90)
where the first and second terms on the right-hand side of (90) represent basal heating and internal heating respectively, and $A/B=2.23$. The inner boundary is isothermal while the mean heat flux at the outer boundary is $r_{o}$. The heterogeneous heat flux imposed at the outer boundary is derived from the seismic shear wave velocity variation in the Earth’s lower mantle (Masters et al., 1996). For composition, a uniform volumetric sink $S_{i}=-1$ is considered, with a constant flux at the ICB and zero flux at the CMB. The basic state compositional gradient is given by
$\frac{\partial C_{0}}{\partial r}=\frac{Sc}{Pm}\frac{S_{i}}{3}\bigg(\frac{r_{o}^{3}}{r^{2}}-r\bigg).$ (91)
By progressively increasing the value of $q^{*}$ at a fixed Rayleigh numbers $Ra^{T}$ and $Ra^{C}$, the relative horizontal buoyancy $Ra_{\ell,H}/Ra_{\ell}^{T}$ at which a dipolar (D) dynamo undergoes the transition to a reversing (R) dynamo is noted (table 5). For a thermal power ratio $f^{T}=11\%$, the polarity transition occurs at relative horizontal buoyancy $Ra_{\ell,H}/Ra_{\ell}^{T}=0.6$ in the dynamo, which compares favourably with the values >0.5>0.5 predicted for vanishing slow MAC waves in the linear magnetoconvection model (table 4). The two-component dynamo model indicates that the value of $Ra_{\ell,H}/Ra_{\ell}^{T}$ at the polarity transition corresponds to $q^{*}=O(10)$.
The evolution of the dipole colatitude $\theta$ in figures 19 (a) & (b) shows the onset of polarity reversals at $q^{*}=20$. In the dipole-dominated regime at $q^{*}=10$, the radial magnetic field at the outer boundary shows well-defined high-latitude magnetic flux lobes symmetrically placed about the equator in the Eastern and Western hemispheres, as in present-day Earth (figure 19 c).
SUPPLEMENTARY MATERIAL
S1     The evolution of a dynamo from a seed magnetic field: The role of slow MAC waves in dipole formation
(c)         B                 C
     Refer to caption Refer to caption
Figure S1: (a) Evolution of the dipole colatitude $\theta$ and the ratio of the root mean square values of the axial magnetic field ($\bar{B}_{10}$) to the the total magnetic field ($\bar{B}$) with time in units of the magnetic diffusion time. The simulation begins from a seed magnetic field. (b) The evolution in time of the fundamental frequencies. The vertical dashed line at point ‘A’ indicates the emergence of slow MAC waves, which occurs when $|\omega_{M}|\approx|\omega_{A}|$, and the dotted line at point ‘C’ indicates the time of dipole formation. Point ‘B’ represents a multipolar state where slow waves are present. (c) The contours of radial magnetic field at the outer boundary is shown for points ‘B’ and ’C’. The dynamo parameters are as follows: Ekman number $E=1.2\times 10^{-5}$, modified Rayleigh number $Ra=5000$, Prandtl number $Pr=1$ and magnetic Prandtl number $Pm=1$. See also Varma and Sreenivasan (2022).
(a)              (b)             (c)
Refer to caption Refer to caption Refer to caption
(d)              (e)             (f)
Refer to caption Refer to caption Refer to caption
Figure S2: Greyscale plots on a horizontal section at $z=0.2$ below the equatorial plane at the instant ‘B’ in figure S1, showing the ratio of the magnitudes of (a) Coriolis, (b) Lorentz, (c) buoyancy, (d) non-dipolar Lorentz, and (e) dipolar Lorentz force terms in the $z$-vorticity equation to the magnitude of the largest force among them. (f) The estimated (lines) and measured group velocities (points) for the blue and red points shown for figure (d). Here, $U_{f,z}$ and $U_{s,z}$ denote the measured axial group velocities of the fast and slow MAC waves, respectively. The dynamo parameters are $E=1.2\times 10^{-5}$, $Ra=5000$, and $Pm=Pr=1$.
(a)                  (b)
Refer to caption Refer to caption
Figure S3: (a) & (b) Contour plots of $\partial u_{z}/\partial t$ at a cylindrical radius $s=1$ for $l\leq l_{E}$ presented for the instant ‘B’ at points ‘1’ and ‘4’ respectively in figure S2(d). The parallel black lines indicate the primary wave travel direction, with their slopes denoting the group velocity. Panel (a) shows the propagation of slow MAC waves, whereas panels (b) show the propagation of fast MAC waves. The dynamo parameters are $E=1.2\times 10^{-5}$, $Ra=5000$, $Pm=Pr=1$.
We consider a convection-driven dynamo operating in a spherical shell, the boundaries of which correspond to the inner core boundary and the core–mantle boundary. The ratio of inner to outer radius is 0.35. Lengths are scaled by the thickness of the spherical shell $L$ and time is scaled by magnetic diffusion time $L^{2}/\eta$. The velocity $\bm{u}$} and magnetic field $\bm{B}$} are scaled by $\eta/L$ and $(2\varOmega\rho\mu\eta)^{1/2}$, respectively. The temperature is scaled by $\beta L$, where $\beta$ is the radial temperature gradient at the outer boundary. The non-dimensional MHD equations for the velocity, magnetic field and temperature are given in equations (3.1)–(3.4), §3. While $\eta$, $\rho$, $\mu$ and $\varOmega$ above are defined in §2, the dimensionless parameters and the basic state buoyancy profile used in the dynamo model are described in §3. The velocity and magnetic fields satisfy the no-slip and electrically insulating conditions respectively at the two boundaries. The inner boundary is isothermal while the outer boundary has constant heat flux.
Figure S1 presents the evolution of the dynamo initiated from a seed magnetic field. The slow MAC waves onset at the point ‘A’, when the Alfvén wave and buoyancy wave frequencies are approximately equal, i.e. $|\omega_{M}|\approx|\omega_{A}|$. Point ‘B’ denotes a multipolar state in which slow waves are present, while the dotted line at point ‘C’ indicates the formation of the axial dipole. Figure S2 shows the relative magnitudes of the axial components of the curled Coriolis, buoyancy, and Lorentz forces at point ‘B’, where the slow MAC wave frequency $\omega_{s}$ is nonzero and yet the dynamo is in a multipolar state. Here, the Coriolis force is balanced by both the buoyancy and Lorentz forces at several points within the volume. The Lorentz force essentially consists of the non-dipolar part of the field since the dipole field has not formed and therefore has a negligible contribution (figures S2 d & e). Figure S3 shows the measurement of wave motion in the dynamo at time ‘B’. Contours of $\partial u_{z}/\partial t$ at a cylindrical radius $s=1$ for the energy-containing range of spherical harmonic degrees $l\leq l_{E}$ are plotted over short time windows during which the ambient magnetic field and wavenumbers remain approximately constant. At the instant ‘B’, both slow and fast MAC waves coexist. The slow MAC waves are generated at several points where the magnetic flux is concentrated so that localized balances between the Lorentz, Coriolis and buoyancy forces exist. The fast waves are present in regions where the magnetic flux is relatively weak. The measured and estimated axial group velocities of the waves at different points within the volume are shown in figure S2 (f). The method of their calculation is given in Varma and Sreenivasan (2022).
The helicity of the slow MAC waves generated from the non-dipolar magnetic flux is essential for formation of the axial dipole (see Varma and Sreenivasan, 2022; Majumder et al., 2024). In the hydrodynamic dynamo at the same parameters where the Lorentz force is zero, the field stays multipolar (Sreenivasan and Kar, 2018), which indicates the essential role of the slow MAC waves in dipole formation.
S2     Mean resultant temperature gradient, $\beta$
(a)                  (b)
Refer to caption Refer to caption
Figure S4: Horizontal section plots of the resultant temperature gradient $\beta$ at $z=0.4$ below the equator for (a) $Ra_{V}=30$ (slightly below onset) and (b) $Ra_{V}=310$ ($Ra_{V}/Ra_{V,c}\approx 10$) at $q^{\star}=18$. A time average of $\approx$ 1 magnetic diffusion time is taken for the supercritical state to eliminate the effect of the perturbations in the convective state. The $Y_{2}^{1}$ heat flux heterogeneity is applied at the outer boundary. The other dynamo parameters are $E=1.2\times 10^{-5}$ and $Pm=Pr=1$.